THG temperature issue, DP2-1. Dawn board replacement not OK.

 Report ppt.


29 May

Problem on Matacq and not on data (not on DQM): instability, for both bleu and green.

Matacq issue: see OM report.

At the end to avoid same problem again (because it was solved by itself ): move cable input from IN1 to IN0 of the Matacq board.

The spare MAtacq was not OK : no signal on lightchecker.


15 May

Reminder of the default setting:

blue: 35A, 40% internal attenuation

new green: 3.5A, 35% internal attenuation.  A is 11,245us for both laser



30 April

Blue DP2-2 not stable.  investigate. during investigation move to DP2-1 few minutes.

faulty DG535 ! (found by measuring the power DP2) . replace with a new one.

send 2 faulty DG535 to E-pool for a quote.


18 March

inversion mentionned after the spybox removal: between EBP03 and EBP04. LM42 and LM43. 2 differents cables, but inversion found on S2G14, EB side (UXCside).

The inversion was compensated on teh spybox sometimes before without notification.



11 March

Clean disk ecal-laser-room-02 (disk almost full).

No new green laser for the weekend : wrong sequence ! old green...


7 March

remove spy box. elog

ECAL readiness review,Marta


20 Feb

reflection test by removing devices one by one  elog


6 Feb

CMSweek presentation of the laser historic, from the begining



defaut laser setting for the run:

new green online:

I=3.5A ; energy=35% (shot102) ; internal attenuator= 50% ; delay A=11.245us


I=35A ; energy=40% ; delay A=11.245us.



reflections  + new EMTC firmware + Adjust WTE by 1 more orbit. elog

Problem of new green laser alignement: OM presentation




new laser settings DBGui : elog

Laser room relocation CMS week


CMS week, laser relocation Marta



Laser room relacation.


test both blue laser + old green, all OK.

touching the EMTC , changing trigger cable => improve the Matacq stability, no more jitter !


DP2-1: 35A, 50% attenuation, A=11.32us

DP2-2: 35A, 50% attenuation, A=11.27ns

legacy green: power=P=250, NimDelay green = 7781




Laser CMS week


August / Oct

New green laser test : report


summary :

1. run at 5A not a 4 (not enought stable)

2. issue with delay using only one DG535: delay is too high so frequency is divided by 2

3. delay of 10ms make the last pulses outside the calibration, in physics.

4. Need to send the WTE 7 orbits before the L1A.

5. pulse was stable with internal trigger on my scope but ext. trigger some problem. Now it is fine.  2 pdf plots from monitoring : dso21473, dso24486


23 June :

move back to legacy system, DP2-2 and LED underground. to take data with B=0 before Monday.

OM meeting, pilot run talk Marta


14  to 25 June : MD  and TS1

problem with laser DP2.1 harmonic. see slides OM  for the summary, with all settings: OM 21 June

last Matacq plot:


12 June:

After HVAC issue : instable temp and RH >40% (spec: 40 max), work onsite to do insulation,ect.

temporary split air unit installed , door leave openned. Temp around 30C inside. should be ready for Wed. the MD !


22-25 May

DP2-1 move upstair. after 2 days OK the chiller start to leak. Change the tank with the one from the spare chiller (no heating).

Laser OK order 2 spares tanks.


PC from IT: 2 pcs . VME control OK after work with Evgueni (problem of viewing the pci board was a problem of privilege. need to be root all the time. slove by jira ticket)

PC for the laser hardware: USB ok but order a new usb to 4 rs232 for the new green laser.


HVAC monitoring: web page Timber from CERN


24-28 April

New green laser installation, Liyuan.

Calibration curve

dso21365-1_4A newgreenlaser.pdf



21 Feb

Laser online for MWRG and stay online all the time with ECAL.


02 Feb

test cable SEDI avec laser DP2-1


16 Jan

switch ON DP2-2 and green for EEPFar check.

Attenuation blue 35%

green power P 240


Ref power green laser Werner:

direct output P(255)=60mW

After 10m fibre and focusing lens: P=70mW



9 Nov

VME PC stuck.  elog.

intervention on Friday with sysadmin. replace the PCI board then the CAEN controler.

After that, laser pulse issue both green and blue: pulse instable in time . visible on Fast monitoring. On DQM is OK because bad pulses are removed.




dso plot, 25days : dso21341-2.pdf

one blue sequence failed: http://cmsonline.cern.ch/cms-elog/1153653


26 April 

no more laser data since yesterday evening: no blue and no green.  Remotely all OK ! => bad DAQ configuration

last plots : dso21397, 13days


14 April

Calibration both lasers : tuning DP2-1 temperature.


8 March 

Switch on lasers. DP2-2 online. All OK

Calibration done just before: DP2-1 degradation compare to begining of LS2


humidity laser 1 : 15%  (25% for laser2) but room humidity Ok, 27%.




15 Nov- 26Nov

Lasers online for the MWGR.

Same laser, DP2-2. 35A, 35% attenuation

Put the spare Matacq board (after reparation) online: Ok for 1-2hrs then nothing.

Put back the previous one, OK. Try again with Evgueni : board not OK.

Get a new board from Saclay (prototype) : test on 25 Nov.


23 Nov: scan test because no data last time with miniDAQ. Run in global this time.

put 99% internal attenuator 11am. 

Laser scan history


11 Nov. 

Laser DP2-2 into error : Temp THG 41.5C . Initialisation not OK.

Replace one Dawn board - interconnected board  : quick test by connecting the cable it was ok.

Ordered 2 new boards as spare.

report/reports2021/DP2-2 board replacement.pptx



14 Oct. 

Lasers ON.

DP2-2 : no reset of the internal attenuation : power too high like it was 99%. perfom a reset on 16 Oct. And put back 35%: OK !



15 July - 13 August CRUZET

Laser ON:

- remove extra fibre TTCci + extra network cable EE

- Delay A= 12.56ns , legacy value.


1) Green laser too low as in June. long diagnostic to find that power is not ok on the laser (adjust NIM delay, etc) : 16mW internal trigger, 1.7mW external.

- Replace laser head with the old one : 50mW internal trigger . 21.5mW after the coupling main fibre + 10m fibre.

= no more spare for the green laser at PT5.

- Power still low => replace the controller with the new laser from Werner : OK ! power back to normal on Matacq.

Keep the old laser head.


reminder : green laser reference direct power output, full power (P=255) : 206 mW, internal mode.


2) 22 July, room temperature : minimum value aournd 11am up to 3pm. run number 343414. real temperature aournd 21.7C

3pm: increase + temp to 5 bars on lcd display. run number 343416


Green: replace laser head with old one + replace controller with NEW one:

Green laser summary

new NIM_DELAY green laser: 7886


7-18  June. MWGR4

Lasers on.

keep the extra TTCci fibre.

No more lightchecker: discover there is no more matacq signal since 15 April.  Try to find the spare

I have done a 200m network cable  : connection on the EMTC side to check the LED : trigger signal 100Hz is there , on USC patch panel LED box. OK !

spy box: can not remove fibres frmo 1x100 switch ! (back and front) will measure the attenuation with laser power meter directly.


17June: get a spare matacq from Saclay. no succeed. Evgueni came to Pt5 and solved the VME issue. elog


DP2-2 : setting 35A and 35% internal attenuation (was 50% perviously)

Green laser : power is very low. don't know why....P=240 as before.


18 June: define the steps for the laser scan , only blue and EE. google sheet

change internal attenuation to 99% to follow the steps for scan..

Blue laser out of range on Matacq while the attenuator is AFTER the pin diode !



12-23 April . MWGR3

Lasers on from 12 to 23.

Green laser: P=240 as before

DP2-2 : 50% internal attenuation. Was 35% before March so cleanning need to be done.

ADC on Matacq: 650

Test: long TTcci fibre of 150m (Tues.13 April):

1. Extra fibre on EB- partition . then connect on EE+, the 4th partition which is actived.

2. not conclusive because it is not the partition-fibre which get the clock. The clock is from the first TTC-rx  . so for us EE-(the top one)

3. 5ns/m so we have +750ns for TCDS signal. Laser light should be 750ns before

4. Change delay A = 12.56ns to A= 11.81us.

5. Green laser: NIM_delay  on ecalpro/.lasercfg : 7901 to 7841 (unit is 12.5ns)  => 750ns ok

6. For Matacq windows: *POSTRIG of laser supervisor : increase by 37. =>  5014 to 5051.  wrong direction

7. *POSTRIG: 5014 to 4977. ok ! blue is back on Matacq. The base unit is 20ns so can not put back the pulse exaclty on the middle on matacq.

pulse  is no more just on time 1490.  Change EMTC_TTC_IN from 0 to 1 (force EE- to be the source of signal)

8. Blue pulse is 10ns shifted on DQM => Adjust Blue delay by fine tuning:  with *LASER_PHASE, from 4 to 2  on EMTC.

[1] https://twiki.cern.ch/twiki/bin/view/CMS/EcalMonitoringControls


9. change the ./monitor delay for fast monitoring: 99120-750= 98370




1-12 March . MWGR2

switch on lasers .

DP2-2 online: internal attenuation 45% give 1300ADC count = saturated APD.

Increase attenation to 50%

issue with ./monitor: pulse no visible or only time to time.




16-20 Nov


PC status: all was configured so PC02 and 05 replaced. Issue with login solved (issue with laser account somewhere else)

Green laser still not OK: no operation of the shutter with the DAQ.

In fact it is because USB port were inverted from July, when VMEPC was replaced.

This  invertion of ports 10 and 11 should be reflected on the configuration file : .lasercfg under ecalpro account

=> The green shutter is controlled by the VMEPC so config. file to update.


PC02: after reboot , the USB port can be swapped also between DP2-1 and DP2-2. Be carreful.





7-10 Oct

MWGR. Switch on the laser only the 2 last days. Wrong GPIB connection before on PC02.

Green laser issue again :

shutter ok with ./shutter program. but not OK online ! 12V power supply replaced anyway.

 solution: I have oppened the shutter with ./shutter THEN online it stayed open because no more control

when green was back : timing out of time ! expected for the last hour ! to be understood.


28 Sept

clean DP2-1 filter => low flow error.


Sept 2020:

All updated program by Evgueni are working fine on new machine PC05.


dp2n, dim, monitor, h4daq.

DNS is starting automaticaly at the booting (crontab)

all program under SAME PC !


Program ok  with root ,  internal trigger : dso-2.ps


work by SysAdmin: https://its.cern.ch/jira/browse/CMSONS-12548


Issue with GReen laser shutter. disconnect and reconnect 12 V power supply help. elog 



15-16-17 July MWGR


Lasers ON. DP2-2 online.

Run with new VME PC and new ecal-laser-room-05 PC .


restart the air conditionning : Vincent Pittin openned the cooling valve.


1) Issue with new VME PC and the CAEN board. Evgueni replaced it by a spare one. 



2) OK for the fast monitoring but not for the ecal-laser-room-02 :  need startech box because of 1U PC but not recognized. the PCIe-PCI adaptor is the same, not recognized to keep the legacy PC for the MWGR.


3) Green laser on VME PC: USB adress is changed !  need to run ./dtl419 10 and no more ./dtl419 11. same for ./shutter.


Doc sumary : https://indico.cern.ch/event/940355/



13 Feb-25 Feb

Lasers ON. DP2-2 online.

Change the attenuator from 35 to 25% on 18 Feb : wrong direction.

change from 35 to 45% (more attenuation)  : ADC from 1200 to 1100.


21-21 January

MWGR: restart laser. New pilars installed on laser room: remove one blue rack and installed one small one.

DP2-2 online:  35A, 35% internal attenation.

Green online: P=240


Matacq : give blue amplitude at 1200ADC : was 1200ADC in Jan.2019 but in July 700ADC .



Laser safety proposal


12-16 august

Laser Maintenance Liyuan. See report

DP2 calibration Plot after maintenance


Quote for DP2 revision


15-19 July

restart laser for MWGR.  DP2-2 online. 

DP2-1 : low flow error on Saturday 20 (after 1 week )



Presentation on possible system upgrade




Long fibre test and EE scan power

Ecal days EE presentation

Long fibre test


15 Jan

restart laser for calibration test: week 3 of LS2. next is Week 10

Green laser P=  240 amplitude.

still laser2. Issue to restart:  on the PC GUI,  current was well 35A but not on laser side -> so no pulse on lightchecker.


Spare optical switch reparation: replace the power supply 5-12V inside. https://fr.rs-online.com/web/p/products/1830111





dismount Quatronix Blue 1. room free from HCAL.


18  July

histogram plot, 400hrs : dso21421-2.ps


18  June

Green laser:  the laser power (controler) is dead. the laser Head is fine.

replace anyway all with new one : new controler and new head.

Retune the coupling. still no ligth on 20June when run restart: problem of too low power. Set as 255 (max P) and pulse is back on light checker.

P=255 -> to high power.

Fine tuning: = P


12  June

Green laser problem, sequence red on light Checker.


1  april

DP2-1: water flow error !. 2.3L/min. Clean pipes, flush all. replace long pipes with shorter one. no effect. replaced the pipe with flow sensor. Back to 2.8l/min . stable. 

DP2-2 still online for begining of physics.


12-16  March

Liyuan visit.

test green laser. replace old chiller DP2-2.

Long test fibre : cannot find the pulse on matacq/DQM. posponed to ..unforeseen date.

the delay on dBGUI LM region seems too high: not possible to put huge value.


8 March

long fibre test posponed : no connectors on one end of the bundle !


21 Feb.

restart laser run. DP2-2 online, 35A


15 Jan.

Laser presenation at the integration forum  https://indico.cern.ch/event/687346/




 13 Nov

DP2-1 flow error. switch to DP2-2 19h00


3 Nov

received new chiller for DP2-2.


14 Sept

Liyuan's visit. Summary here.  New chiller ordered.  the black on on DP2-2: still leaky. The spare one is the old one which need the separate heater device.



15 August

blue OFF DP2-1 : Guy oncall. LED shutter off, laser into error. current was well 35A) reset the laser itself. then ok .elog


18 July

after another water flow, no more water (leak) on chiller : replacement by a spare one. (old one from laser1)


11 July

Water flow error DP2-2. Swap with DP2-1 , 15h40.


10-16 June

Issue with laser stability on lightchecker  : http://cmsonline.cern.ch/cms-elog/990383

problem : connection of the batterie and pin diode Matacq! soldering again. need better fixation .



receive new DG535 (ordered from CH)

received 2 flow meters as spare + one Laptop


5-9 April  

Liyuan at CERN


8 March

Laser 2 DP2-2 online following low flow error the weekend on Dp2-1 --> new pipe with new sensor, since OK then. Flow rate is 3.3/3.4 l.min

General power cut on 9 March: only lasers went OFF, not control room. Green laser restart with internal trigger by default.



restart lasers: maintenance DP2-1 : flush cooling pipes because of low flow rate bellow the limit (2.6 l/mn). now it is 3.2l/mn.

clean filter. check attenuation of the optical path (1x100 switch). clean fibre , OK

Green: put back external mode.

PC: ecal-laser-room-02:  connection is bailleux@cmsusrpup only ,  not .cern.ch ! reconfigured GPIB ok .

Problem because no more .dat file on slow monitoring --> on ecal-laser-room-02: #printenv  there is no DIM_DNS_NODE ecal-laser-room-05 running !

need ot configure again with : ecal-laser-room-02# tcsh, then #setenv DIM_DNS_NODE ecal-laser-room-05

--> start again ./run on dp2n, ok !





26 Nov

saturday . bad DQM data: laser too hig on EB, too low on EE..  elog. eems wrong attenuation.  Go to Pt5 to perfom a hard reset of the SHOT102.

tehre was no error on log file, attenuation well moving from software but was wrong information: no way to be detected other then d<ata quality on DQM.

(remote attenuation is after the matacq or fast monitor so cannot be seen).

A suspect faulty SHOT102 after touching the DG535 which is on the same loop.


25 Nov

new DG535 for DP2-2 : no more jitter, back to delay A=12.560us. --> bad DG535 box as in Feb2015.

Keep DP2-2 online


23 Nov

degradation of DP2-2 stability: jitter and timing. Plot here

switch to Dp2-1 at 1am


18 Nov

Notify an increase of humidity on laser room by a factor 2 in one week ! plot here


9 Nov

error flow again on DP2-1 : clean - flush the pipes and flow back to 3.0 - 3.1L/min. Was down to 2.8 before.


9 Nov

start green NRJ scan . elog here.

last scan was done last year :

green : http://cmsonline.cern.ch/cms-elog/893110

blue: http://cmsonline.cern.ch/cms-elog/903371


4 Nov

low flow error on laser: fake alarm! all ok when turn OFF and ON.   B field was finished to ramp up at around 6pm. laser error at 8.30pm.

Change to DP2-2, spare one from Friday 11pm : 35A and 35% internal attenuation.


3 Nov

restart lasers, PC. and LMFC on ecalod machine.


1 Nov

Stop laser data taking, magnet ramp down. all OFF as (PCs off)  because of 400KV intervention.

lasers DP2-1 was 35A and 30% internal attenuation.

fast monitoring was ./monitor 99120


25 Oct

Liyuan visit. Quick check of DP2-2, only calibration. No other things because of data taking.


7 August

Low flow error from chiller spare DP2 , DP2-2.

16 august after vacation: a water pipe leaks.. under the thermal foam.  Cannot open fully the chiller , need to cut the pipe and add extension.




Swap laser PC supervisor with spare :

. put CAEN PCI board on it : from first PC, or from Evgueni or from sysadmin.

. request changing alias via sysadmin

. change pc connection

. reboot vme crate - clean fibre if not working.



11 May

laser supervisor PC reboot and not possible to restart the run elog

no conclusion why at the end.

following that try to see how to change with the spare PC : vme-pc-laser-room-01 to 02.

first attempt :  not so easy and no time because of beam... borrow the CAEN PCI from M.Dobson.

changing Alias of the PC take time.  need to validate the USB green laser as well.

Next try  during TS1.


5 May

plug a new USB sensor key on vmepc-ecal-laser-01. temp plot on lightchecker for Marc.


28 Avril:

B field, 3.8T



laser room temp : need again for Marc.. see here DPG


- average value no more OK . 12deg for the 4 probes.

- the one display on the DCS in blue and green are data from our laser DAQ, not from ELMB.


problem solved restarting the ELMB :  Jira DCS issue


plot of all ELMB temp sensor can be view not on DCS panel but on the cmsonline Generic plotter:

choose ECAL, ECAL laser room temp.


2 Avril:

Dp2-1 power a bit lower than before while should be higher : best coupling and main fibre polished. pin diode to be retuned.

same calibration as in Feb.

laser3: idem.


29-1 March: liyuan.

1) The output power of the two DP2-447 lasers is measured as function of the current of pump diodes, no evident degradation is observed. The 1st DP2-447, which has been online since last year’s maintenance, now has a total runtime of 15K hrs (about 2 times of the warranty time), indicating that running the laser at 35 A pump current (with a maximum of 50 A) increases the lift time of the pump diodes. The pulse energy, width and jitter are evaluated with internal trigger and a short term run of 1~2 hrs, they are all well within the specification.

2) The ancillary optics of the two DP2-447 lasers are checked and the total transmittance is measured, no damage and degradation are observed. The output fiber of the 1st DP2-447 laser is polished and cleaned, and the coupling alignment is optimized. While the power measured after the 1x100 optical switch is increased by about 10%, the monitoring pulse amplitude at the fast monitor and LightChecker doesn’t linearly increase, which may be caused by mismatch of the area of PIN diode with the beam size or misalignment of the beam on the PIN diode in the monitor box.  More diagnostic is needed to figure it out.  If it’s true, we may insert a beam matching lens in front of the PIN diode to improve this power monitoring linearity in near future.

3) The Quantronix Ti:S IR laser is also checked and tested, it’s basically in working condition, but long term stability may be a concern. If it’s needed, a DPSS Ti:S laser with affordable cost would be an upgrade option.



calibration of the 2 DP2s

16Feb: lasers online, green and Dp2-1 : 35A, 30% internal attenuation.




read Vinj value: ecal-laser-room-05/home/laser/Vinj# ./Vinj

slow data Sept_Dec.jpg       Slow data Sept_Dec_temp.jpg



calibrateur  Vinj and yoko7555 : try to connect back on PC to get the status online.



Technical stop , no laser but found the green laser running !

in fact it was in internal trigger mode !

the control of ./dtl419 11 was not responding : can't control the trigger, frequency...

The 2 USB ports was reverse ! why ?.. ./dtl419 10 instead of 11...

--> changed the connection of the 2 USB ports.


25 Oct:

Vinj stop. Seen by DQM test pulse. (just same day as B field ramp...)

Same thing 10days ago.




In oct, 2-3 hard reset of VME PC . unknown reason.

last one : CPU failed on Friday 6Nov.


1 Sept

calibration, all ok.


20 August



6 July

B field = 3.8T


15 June

TS1 technical stop.

maintenance cooling ventilation laser room.

green laser: ok . see on last plots


11 June

Green laser back online:

- delay adjust in side laser supervisor file : ~ecalpro/.lasercfg -> "nimdelay = 7887"  or 1ECF.

it was 7897 originaly. then I found good value 1ECE=7885 but a bit toow low: based time is 12.5ns so  7887-7885 = 2x12.5=25ns.

not needed to change the firmware (it was changed anyway on the spare EMTC)

- green_laser_power=5 % on dBGUI instead of 20. for the EE it is : 5x8=40%.

- new shutter installed.


29 May

Green laser:

tuning delay register 28 EMTC: move from 1ed9 to 1ecd to get the good timing (7897 to 7885)

this value is reset after each new run so need to fine a solution.

amplitude: changed  internal power to 160 , which move the pulse as well in the good direction (matacq timing)


13 May

Alu piece for the green laser done.

re-tune fibre mounting.

green power : 215mW.

after 10m fibre + coupling to the main fibre: 52mW: 6dB.

the delay should be the same as the old laser but nothing on Matacq for the 1st attempt.



5 May

Remove IR from the sequence. (calibration time to time because of first collisions..)

IR data to be analysed...


23 April

put IR laser on the sequence.

install new electronics for spy box.

laser data from DIM server ok , PVSS up to date now.


17 April

finish to put back IR laser.

calibration plot ok. new lamp connector because of corrosion...

ready for Monday if needed.


Laser system description review (on technical doc.)


10 April

- calibration were stop during splashes events , do not lost data of physics.

calibration restart now.

- proceed to the EDH order for spare parts and consumable. see technical doc.

wait next year for diodes.

- test the new PCi expansion box from Gigi. all ok under B field.

- green laser should arrive 11 May.

- replacement in progress of  old spare PC ecal-laser-room-09:  should be vmepc-ecal-laser-02.

- need probably to put back RED laser.


27 March

- intermitent power on laser dp2-1  on matacq :  shutter error on logfile.

restart the soft solved the issue...  Matacq plot.


- new script to get a logfile automatically: ./run instead or ./laser.

- fast acqiris: remove line 247-250  in monitor.c  to remove time out text on terminal and see the trigger rate.


23 March

. dp2-2 with lower attenuation for Gain 200 since few days.

. order send for a new Green laser

. problem time to time loosing some events : less than 600 ! so no data on some 1/2 SM..

 -->  new plot on lightChecker by Marc. to view if all SM got well 600 events.

fast acqiris plot 8 March: can see data to 0ADC.

need to run logfile to get the time of optical switch.

. add reset SHOT102 ok for Dp2 : test prog. under mydp2 folder.


17 March

stop run for B field.

Matacq 16March.pdf


6 March

green laser: shutter is ok but laser problem now.

elog here


4 March

green laser : shutter problem, which explain the on/off on the sequence.

after investigation, shutter no more reliable.  lost green , shutter block. only dp2-1 !

need to remove the green filter (wheel) which is damaged by the beam...


3 March

- plotting green laser on fast : show long queue for the pulse timing and 2 mains peaks... EMTC trigger ? or new Acqiris ?

why not red dots on matacq ?

- lot of issue with calibration&EG trigger .

- ask to remove DP2-2 from the sequence.  Main laser : DP2-1 for Saclay (old one)


26 Fev

The  pulse position of the DP2-2 changed  for unknown reason 

check with trigger in local and pulse well moved by +40ns compared to Dp2-1

need to change delay A to compensate...

moved  A=12,575us to A =12,545us.


24 Fev

fdesign to modify the laser GUI : laser_gui.c



water flow error  Dp2-1 around 0am. water flow too low at 2.5l/mn.

just restarting the chiller and laser : 2.8l/mn , error no more there.

Flushing the laser head help: 3.4l/mn.

goes down to 3.1 after few minutes



change monitor full scale range during run: 2v -> 0.3V for higher pulse..

the 2 DP2 are now with same level :  matacq_17Fev2015.pdf

after polishing main fibre on 9Fev for both laser, the DP2-2 is going done from 1300 to 500ADC... ?!

suspect linear  attenuator problem.. because both laser have quite same output level in mW.


dp2-1: before coupling, 38.0mW 25A, after 1m coupling 33.0mW, 100% attenuator

dp2-2: before coupling, 47.7mW 25A, after 1m coupling 41.3mW



12 Fev

start Cruzet


9 Fev

Check power of DP2s. calibration done after attenuator. check linear attenuator and both ok, 100%=full power.

polish main fibre of both laser.

power output DP2: DP2-1, 90mW@35A, DP2-2, 85mW@35A

DP2-1 should be ok on matacq now.


3 Fev

readiness review : one DP2 online is ok now, cf. MArc.


29 Jan

received trigger in continous mode but stack on green laser ch. 88 . see elog 838065


- readjusted delay because of new monitoring pin diode for lightChecker:

DP2-1 m delay A=12.74 us, (was 12,75)

DP2-2 delay A= 12,575 us  (was 12,58)

- power too low for DP2-1 ? was 1050 ADC before, now it is 450 as for the green.


28 Jan

MWGR: News for the Lightchecker:

form Marc:

Comme pour l'analyse des fichiers lmf, on utilise la pin connectée au canal 1 du matacq, je propose d'utiliser la même pour le lightchecker.

the pin in used for the lightchecker changed from 0 to 1 so lightchecker data  may be different  !



restart all after 1 week  because of no network at CMS.

changed the patterz for the spz box.

found opposite direction for the 2 DP2  internal attenuations

0% for DP2-1  mean 100% of power,

100% for DP2-2 mean 100% of power.

Keep the same config as before, 29% and 35% for DP2-1 and 2.



modification of the safety box ok: shutter interlock for DP2-1&2.


15 Jan

no data on EE-3 -4 -5 : problem was on spy box, channel e-f-g-h not in same order as the others so wrong labelling during installation

swapped fibres 86 and 87. 


reparation of the SHOT102. new Capa. should be ok.


13-14 Jan

laser run with new monitoring boxes installed.

EB- and EE only. same attenuation as last year to compare.

seems amplitude ok on DQM: 1037 min, 1649 max

seems no data on EE-3,4,5

try to restart the Matacq but no data (start lmfc ok)


6 Jan




15-18 Dec

Installation of the spy boxes + test output with lasers. JLouis&Inna.

- one shot102 failed: burnt. no more working.

- dg535 of dp2-2 replaced but to be checked again: light ok but no more display.

- output power ch1 spyBox: 5mW . was 4.2mW before (1x100 switch) excel file


11 Dec

Adjusting intensity of the DP2s:

1) normal value :  DP2-1@29%, DP2-2@35%, both at 35 A --> to get 500 ADC counts

2) highest value for central EE :  DP2-1@52%, DP2-2@63%, both at 35 A --> to get 900 ADC counts



9 Dec

DP2-1 : 35A, 35%

DP2-2: 35A, 42%

running on new script #dp2n to get 2 DP2 online.

-> SHOT102 to be reseted after power cut or long interrution else lost position ! (no light on ECAL)


measurement of the power at the channel nb 99@35A and attenuator at 99%

DP2-1: 5.1mW

DP2-2: 5.2mW


27 Nov

updated of the laser code to have 2 photonics running online: new folder is #dp2n

0, 440-TiS

1, green

2, DP2-1

3, 800

4, DP2-2

see presentation last reports


13 Nov

update slow DP2-2 histogram + matacq as ref. from the 12Nov with 3.8T.


12 Nov

laser setting on DB Gui: http://srv-s2f19-30-01.cms:5000/static/index.html#/login

TTCci config -> view the sequence, blue, IR...

laser config -> view the attenuation setting. bellow: xml clob : edit -> view the setting for each LM


in order do not have DP2 fired and laser data when Quantronix is on the DAQ sequence, while laser is OFF: change IR laser from offline to online ?!


11 Nov

global run with B field:

DP2-2: 35A, 100% internal attenuation. was 3% before.

from elog: DQM is ok , ~2000adc for the pulse on DQM (was 400 before)

Ask to reduce power ~25% for the DP2: on Matacq it is 1400ADC so close to the limit ! should be 1000 on the Matacq.


Looking myself at the DQM there is saturation for almost all EB so reduction justified.


calibration sequence view from laser daq:

DP2-2: remote attenuation : 9% for EB,  81% EE  (internal = 100%) but as well 5% for EB and 81%EE.. changing time to time... !

green: remote attenuation : 20% for EB,  99% EE


--> understand problem on Matacq-config: the Quantronix was still on the sequence ! so when calling the Quantronix, the DP2-2 was still selected,

because QUantronix offline, but with settings from DB Gui of the quantronix: different attenuation !


--> settting of the DP2-2 to not saturate : 35A, 50% internal attenuation.


7 Nov

check attenuation on the DP2-2 chain:

first measurement give almost 0mW on the channel 1.

after a reset of attenuator, attenuation ok: 12 dB on ch1.

Can explain the low power on the APD (factor 2)see by Saclay-DQM.


14 Oct

switch the 2 chillers between DP2-1 & DP2-2:

DP2-1: still 25deg; DP2-2: 26 deg. can't put 27.8 deg with this old chiller (no heating capacity)


9 Oct

DP2-1: without the heater part, no effect. refill water after 3 days (~400mL)

- MWGR8: ask to change the Matacq legend ; and to remove IR laser. seems low signal on APD now x2 ?!

- vmepc-laser-01: disk full can't login anymore. call DAQ oncall. elog 825229


6 Oct

DP2-1: no more water after ~3-4 days.. (water filter in good direction) water flow 3.6L/min

Remove the heater to see (goes to 3.9L/min)


29 Sept

Run with external trigger (not calibration sequence) DP2-1  and DP2-2. SHG temp is stable at 0.0deg as it should be ! solved !


DP2-1 water consumption: Found wrong orientation  for the filter: the arrow si in opposite direction of the water flow. Inverse the pipe.

The water flow increase imediatley from 3.2 to 3.6L/min ! maybe solved the water 'evaporation' issue


26 Sept

DP2-1 Fault temp solved. open laser cover and touched LV connector.

See elog 824256


22 Sept

DP2-1: laser Fault after 3-5min each start :  " SHG temp error".  Temp are unstable +-2.5deg while it was +-0.0deg at begining.

clean water tank , filter. some dust anormal.

in touch with Photonics.



DP2-2: clean filter


1 Sept

With JM Andre@pt5 (sysAdmin):

PC05 was not starting properly because of wrong modification with VNCserver by him (vncserver modified to start automatically but can not)

VNC is removed and just using X11vnc it's working fine

-> external power 12V for PCie is not supported by Dell...


 --> put PC05 on the rack with Acqiris, all is Fine !


1. Change file ./tcshrc on PC02 to get good dns server PC: PC05 (was PC01)

2. #./ monitor 99150 (was ./monitor 99000 on PC01)

3. Start dns on PC05 with

# home/laser/dim/linux~$ dns

4. # ./h4daq



PC05: still no graphic mode on boot but ok with startx  and ssh ...

found on ATX connector the +12 V is there but connector is faulty time to time thus only 5V and no 12V on Acqiris.

Acqiris is OK:   pc05_monitor# ./monitor 99150 pulse is there !


End of WGR5 this evening but no laser data.



IR laser: After investigation, found the  logic Unit to be faulty. replace the one from TiS1 and current is OK. Calibration done OK.

Fast monitoring ok on PC05 but trigger was ok before and no NOT ! also problem on PC: power12V not here? no graphic mode,ect. !

problem with power supply on PC02: no power from the power cable: fast Acqiris not power on.

PC05: Jira ticket: no more graphic mode..



IR laser: try to tune it, doing service after MWG4. Found that current is no more good: always YLF=1080mW whatever the current is: 7A or 20A !



Last Matacq plot: DP2-2, IR, green. Error on display: there is only one blue ,  IR=DP2-2,  Blue= TiS-blue but it is not online , should be removed.

Reduced IR delay A.


DP2-2 still @ 35A.


in April: Dp2-2 amplitude was 600 @40A 50%, now it's 1000@ 35A/3%   !! 

--> maybe due to not the same optical switch 1x5 connector



Lasers online data taking with miniDAq. no pulse on matacq but pulses ok they are here. see elog

- histograms online of DP2s lasers.

- Power too high. may result of the cleanning campain of all optical connectors.

-> change attenuation from 50% to 20% for DP2. then Changed again to 3%.

-> changed current from 40 to 35A because of no effect of attenuator bellow 5%.

 leave 5% for IR


23 July

- new 220V power bar ok

- new VMEPC finally working with mini daq today (new driver VME ok now..). see pulses on xdaq monitoringm not yet on Matacq (nfs problem on pc supervisor)

- config file on ecal-laser-room-02 in order to keep the USB and GPIB mode r+w all the time: /etc/udev/rules.d/


- list of linux commands for new PC VME. (installation,ect.)


15 July

- Nuno Dos Santos for adding a power bar 220V.

- EDH to ship quantronix laser2 + parts.

- internal run DP2-1, dso21375-1.ps still the unstability of THG SHG.


14 July

internal run DP2-2, dso21370-2.ps 

run with external trigger to get and to analyse fast monitoring: 20140714-0.hbook

no good ps file and error when DAQ stopped but .ps are OK when processing the .dat file.


- DP2-1: no water consumption after 5 days with 0A. Try again with 40A.. no difference after 1 day. wait and see..


- new laser supervisor PC for ecalod-laser-xdaq:  pc name is vmepc-ecal-laser-01, alias is ecal-laser-room-01 and ecalod-laser-xdaq.

- removed ecal-laser-room-04.


2-4 July

global run. new rack PC ready . re-installation from scratch ok after several days (to change the machine name. )

DP2-2 online, 40A, 50%.

will change to DP2-1 later if needed.

IR laser online.

Dns server : keep running on ecal-laser-room-01 but could changed with ecal-laser-room-02. to be tested with ecal pvss,ect.


- CMS IT bring new rack PC for laser supervisor. removed spare ecal-laser-room-04

- DP2-1: anormal water consumption. discovered since spring already. after 2 days of data taking 21-22 June, water flow error-> no enought water !

test only chiller+heater over 3 days (28-30June): no  lake of water at all !



new DP2, DP2-2 online for EE. put fibre and cable instead of DP2-1.

DP2-2: 40A, 50% attenuation, A=12.59us.

DP2-1 was: 30A, 70% attenuation, A= 12.68us



4 lasers online on demand : IR, TiS blue, green, DP2-1.

TiS blue working the 10April (no light before) and IR power was too hight the 9April (no time to do good tuning. short notice..)

Red laser: 20A, 5% attenuation. TiS blue: 23A, 99% attenuation

     elog 816878     ;      Matacq_10April_2.pdf        ;        CMS_wedn_meeting_WGM.pdf       




Global run. keeping old configuration (1st DP2 with ecal-laser-room-08,ect...)

Change delay A DP2-1 to match the good timing on matacq: A= 12.68 -> A=12.735. (don't know what changed..)



second DP2 installation.

clean and polish attenuation box fibres. Maybe the power wil changed for 1st DP2.

elog 816522



online run, dso21791



DP2-1: temperature study (plot on histo 2014)

SHG&THG were stable up to January (internal run 21762, 23 to 31Jan) then Feb not stable:

+/- 0.3deg. run 21764 14 Feb. 

-> No effect on pulse energy -> histo OK can't explain the double peak with external trigger.

-> Room temp. no impact (22.5 or 24 now) diode temp. always ok but not THS-SHG. (chiller doesn't changed)

-> current at 0A or 51A doesn't changed the temperature as well.

-> new SHG&THG settings: no effect.


Remain: settings of temp SHG&THG unchanged : 49.8 and 50.8 DegC



re start laser run EB+ for weekend Matacq_DP2-GREEN-14March.pdf

Fast monitoring -> unusable data 20140316-0.hbook. Wavelenght WL wrong ? because got TiS blue data ! and NRJ green or DP2 wrong...

slow DP2: dso21781. Temperature diode-shg-thg unstable  !!  -> why only with external trigger ?



plan for 2nd DP2 installation: 2-8 April.




conclusion double peak:

- marc remove already the first 10 samples of each bursts so no problem;

- same effect on 2nd DP2 but lower effect;

- photonics: effect can be minimized by adjusting the SHG-THG temperature.



Turn lasers ON for data taking. re-start Matacq process on ecalod but PC03 not reacheable friday (network cable)

intervention Guy.

run 2hrs DSO21764. investigation to understand the 2 peak in amplitude -> the first event is lower , see with a DG535 external trigger

-> found that the 2 peak was there in Grin, Nov2013. first online run after receiving laser from US.

before reparation, run online  in Jan-Feb: no double peak ! -> new laser hardware (xcal,diode)  ? 



DP2 : measurement done on 1st laser to redo the performance%current . See calibration file.



DP2+ green for EE --> data taking ok

DP2: 70% internal attenuation @ 30A.

Lighthecker not working: found command on twiki lmfc start/stop/status matacq_feedback  restart seems ok ..



Restart DP2+ green for EE but no data taking.

DP2: 70% internal attenuation @ 30A





start DP2+ green for EE, short run.

DP2: 70% internal attenuation @ 30A


8 Nov

IR laser not working, lamp, GPIB no control.

 current= 20A, attenation=10%.  pulse not stable : TiS Qswitch tune several times..


7 Nov

IR Laser: start. new lamp.


5 Nov

remote attenuation for DP2:

45% for EE; 5% for first channels of EB 


4 Nov

GRIN; global Run In Nov. DP2 online.


current = 30A;     delay A=12.68 (was 12.7ns first data)     Attenuation= 70% (arbitrary).

remote attenuation = 45% for EE Ch 81-88 (99% for the Green)

Red IR online why?

Matacq plots


removing the IR from the sequence ~4pm



Quatronix+green for EB- test.


delay A= 95.54us; Current 23A; internal attenuation = 90%  (before 100% because low pulse)

on Matacq , amplitude is 700; 820 for the green.

1st DP2 ready for GRIN. all OK. change SHOT102.c for new attenuator.

clean main fibre of DP2 (switch) , Quantronix and green.


internal run at 30A ok . good result.


11 Oct

Get 1st DP2 at CERN. Bring it to PT5.  First swich on Monday failed: low flow error. 2.2L/min.

change the water pipe from parrallel to serial link -> 2.2 to 4.3 l/min. ok. Can't change the level over the software (not the same version)

Spare chiller (the new one) not working.

DG535 get with Photonics: never able to switch it ON. fuse brake each time.


14-15 Oct

short run with quantronix laser1 + GREEN  for EE only.

power goes down moday afternoon after room temp. changed : 22mW@23A. after tuning the 15Oct: 80mW.


11-12 Sept

short run with quantronix laser1 + GREEN  for EE only  data here:  dso21495-1.pdf


15-16 July

short run with quantronix laser1 for EE only (no LV for EB)  data here:  dso21489-1.pdf

Like the green shutter not working ..



Connection of  the Startech expansion box on ecal-laser-room-04.



short run with quantronix laser1. for EE only (no LV for EB)


end March

training 3 days,  Photonics. see on reports

DP2-1 arived at Phtonics for reparation.


19 Feb

switch OFF all . DP2 laser dismounted and on transport the 21Fev.  Parts for Caltech alos shipped.

training at Photonics foreseen 11-14 March.


18 Feb

optical chain power measurement . see report here.

green laser scan from weekend. bypass shutter (prog. laser from Philippe). why? no reaosn becasue shutter was ok again . safety optical switch..?


13 Feb

green laser : OK again ; changed shielding orientation and also a screw was unscrew and the shutter touch it each time. elog.


7-8 Feb

green laser : shutter BAD again . see elog


4 Feb

green laser: installed MU shielding around laser -> shutter OK !


30 Jan

green laser back to normal without any intervention !

before shutter stay closed half of the time. mystery.

after 1-2 days, back to fail again...


24 Jan

Laser1: it's recovered. found the problem of the flow switch error.

touch the green laser shutter to debug it: it was closed half of the time ! suspect also the b field..



20-21 Jan

Laser3 offline on saturday because of out range; tuning the 21 Jan and put online the afternoon.

Laser1: new logic unit, sensor flow, valve and power distribution: still water flow error !

15 Jan

Laser3 - IR laser online . Laser1 was anyway stopped because of too low power -> YLF crystal damage, water inside !

14 Jan

restart laser1 after power glitch on weekend.


tune Laser1; remove TiS Qswitch,ect...


Switch ON lasers

Laser1; new lamp. power too low.




Going back during holidays to switch on back the 220V for the DP2 !!!!


17 Dec

Switch OFF lasers by Ken...


12 Dec

- get the spare Acqiris board

- DP2 current raise from 45.2 to 45.3A


29 Nov

DP2: change feedback setting to original one: 700 ADC counts and not 400.  current was oscillate 0.1A with 400 count..


26 Nov

clean main fibre DP2 side. calibration of DP2 (changing current).

B=0T Monday 26 ~10am.

B=3.8T 28Nov 17h.

19 Nov

restarted PC 02 because sreen freeze... work on laser supervisor. change it with spare PC-09 later on.

when restart DP2 program, feedback was 400 ADC while it was 700 before (on the log file )


19 Oct

Laser1: reach 23.5A with feedback. changed it to 23.0A and adjust the delay A=95.42us to keep the same timing.


2 Oct

Laser1: no more light. lamp out. replace the lamp, new cartbridge filter. small tuning: power lower than before. run at 22.5A.


26 Sept

Virtual OM meeting. laser plan


17-21 Sept

Liyuan.  Actions for DP2 (investigations,ect..) B=0T.  elog summary 785751


13 Sept

Power measurement

replace chiller with new one: to be ready for next TS, to add the heater to this chiller.


3 Sept

Power measurement. Remove shielding . Power: -14% from 2 August, unkwon raison.

Google doc : report regarding Photonics laser. Link on left hand side menu.


28 August

11am to 15:00 : no DP2 laser --> broken fuse on chiller DP2. (step on width at 10%)


24 August

Installed shielding on Laser head


21 August 14h to 22 August 23hrs: B field OFF  14h elog 781555


20 August

Laser1 out: new lamp, new flow tube.


16 August:

laser PC08 rebooted : USB error. sr470-1 is disconnected -1

--> You need to run "chmod o+rw /dev/ttyUSB0" with root privilege.

removed HP-1 because seems there is error.. no SLOW data from 16 to 20 !!


11 to 14 August

B field OFF 8am to 14 august morning.


2 August:

DP2 power measurement. beam dump at 12h.   See report PDF file.

Spot on IR filter close to the DP2 output. I didn’t want to touch fibre to understand clearly all actions.  Also I have limited time, beam was imminent.

have tested ~10sec the DP2 with a trigger (put trigger source instead of EMTC)  to see our fast monitoring answer, and power level  is ~15-16 ADC,  while it was 13.5 as you see on my previous fast monitoring data. 15-16 ADC count is like after TS end of June,   thus I don’t expect saturation on ECAL.


11 July:

DP2 power measurement. See report PDF file.


3 July:

water leak air unit..elog


2 July:

Laser2 online Sunday ~9am because laser1: no more pulse: only the YLF shutter was closed...   elog 774012

Put back laser1 during interfill at 14h40 the 2 July.

DP2 failed also the SUNDAY: shutter error ...

DP2 prog. crahes : when pressing exit.. no hbook files.  


25-29 June

Technical Stop. Liyuan.


Magnet OFF from 19June 6am  to 28 June evening 5pm.

- DP2: 2 news mirrors on DP2. +2 degrees laser room.

- Laser2: new crystal, lamp and YLF LBO

- Fast shutter on laser1 and laser2, DP2 , all with USB.

- Laser1: new lamp.


12 June

after new chiller only half of the DP2 power on the Matacq !

17h10: beam dump time to investagte. See report (OM on 11 June) SHOT102 problem (internal attenuation). Clean also fibre.


11 June

 beam interruption --> replacement of the DP2 chiller on the morning.

- check the power before replacement of the chiller.

- filter on the water pipe was very dirty. clean it.  Put 1L of water inside new chiller + 0.5 ml of each solutions.

- set the temp . Ts= 25 degree at 11h13. At 14h25 Ta= 24.4 degrees.... !


6 June

Shutter error DP2 : when going back to Global run after minidaq  elog . See Matacq plost on histo2012. 

No intervention ; shutter back to normal after 2hrs of global run, maybe due to new script laser1 to reopne the gpib connection automatically.


4 June

- DP2 still going down: ~700 ADC bellow the green.

- no more GPIB error

- laser1 still OK

- Put Fluke temp. recordered.


20 May

Sunday 5:35 am cms electrical glitch. Laser1 stop. get call from pt5.. transfer to Jan V. restart laser1 ok ~13hr.


16 May

DP2 447 goes down and down: was 1000 ADC now it's 700.

Thus not due to new room temperature following friday. elog to request investigation.


15 May

laser sequence stuck , 1 red sequence when we was in pt5.




friday 11 Cegelec intervention on the air conditionning. only 1 controller now for the 2 units. Set point 20 degree when set to minimum temperature.


9 May

DP2-447: gpib error at 4:25am. discovered myself at 7:20am. Solved it ony by sending command GPIB_CONFIG -minor...

nothing to do with ICS box or hardware directly. Suspect bad driver,ect.


8 May

Green laser out of sequence for 1 sequence. Guy's investigation but  crashed of its VNC and also ecal-laser-room-02 log out !!!

--> laser DAQ stopped... error for ECAL...

restart by Emanuele underground but lot of issue or problem to restart DP2, ect.

(DSO laser2 adress change to 7, ect.)


7 May

laser1:  put feedback ON. small laser DAQ intervention...

remember lasers setting : internal attenuation = 40% for DP2 and Laser1.


2 May

Laser2 online by Emanuele during the night.  laser1 OFF 'water temp omega' ?..

power glitch 400KV CMS.. maybe part of the reason...


investigation: laser1 : water flow error. list of actions:

--> reset power supply, swap the 2 relay chips on logic unit,  replace the sensor inside tank :  didn't help

--> replace the right hand side chip relay solve the problem.

Restart laser1 : lamp was broken...

remove piece of glass OK without dismount the flow tube.

same power as last time 13 April thus should be have no difference on monitoring.

power = 20.5A  (because of 20.5A was put the weekend by Emanuele)


power LASER1 = 20A after 2 May (too much power) 40% attenuation


email summary here  (OM list)


1 May :

Need to reboot the ICS box Photonics by Emanuele during the evening-night:  shutter error. (same times it was a run with lineraty scan with DP2, strange)


24-26 April

Power measurement done during technical stop.  See slide OM meeting here.


23 April:

If ecal-laser-room-08 is rebooted: run this command under root (su -) : gpib_config –minor 0 (& 1 for the second PCI board)

See automatic problem with Emanuele while rebooting PC. Confirmed by Kejun.


18 April:

Change current DP2 from 45 to 55A (to get 800ADC, was 450 before) Delay adjusted .


16 April:

DG535 spare for the green laser. before using timing delay through EMTC.

DP2-447 manual online


6-13  April

After 1st data taking in global run: lasers several issues for lasers from Easter.

- laser1 unstable: failed to put laser2 online the sunday..keep it all the week at 25A because of Photonics priority and Marc decision.

- Photonics: shutter problem due to GPIB. Wrong PC name as well monday.. no communication : Pedro changed the name of PC

and-->  Replace the ICS box with the laser3 one. OK then !


Laser1 service Friday 13: new lamp, remove flow tube because of soldering needed. All ok after at 20A !


19-29 March

Liyuan& Kejun for laser installation.

new configuration:

ecal-laser-room-01: unchange. Need to start always Dns server on this machine.

ecal-laser-room-02:  run monitoring  on /home/laser/laser# ./laser

ecal-laser-room-08:  /home/laser/dp2-447#  ./laser

ecal-laser-room-08:  /home/laser/dp2-447#  ./avs for the spectrometer if needed.

test DAQ communication in the same way: ./client ecal-laser-room-02


New  DP2-447 laser : configuration file is the same, laser_h.cfg but H4DAQ need to be disable, useless. communication is done throught internet by calling this PC from ecal-laser-room-02.

On GUI interface:

LDD ON:  means the diode are ON (laser ON) 

Qswitch is always ON.

TDiode is the temperature of the diode

Iset is the working current, 45A now. Increasing or decreasing the current to turn ON laser or to turn it OFF is done by automatical step --> 1A/ sec.

Delay A: main delay to be adjusted if needed to get good timing. Others delay should not be touched.


ecal-laser-room-04: green laser DLT-419QT from Saclay. No visible interface.  Manual here.


Work on Quantronix laser:

Laser2: new TiS LBO by Liyuan . no effect. tuning to get FWHM smaller.

Laser1: water leak on YLF LBO. replace the LBO . tune laser1 YLF and TiS myself.

Laser3: no tuning, OK.


13 - 16 March

Marc&JL: green laser installation.

first test online : channel 5 optical switch, program instead of laser3 . plot here.

Laser1 online as main laser.




finishing electricity job.. Laser1 : not  put back on: no air conditioning, wait installation of controller..

Received SIMPLEX cleanroom. Next day the optical table.

CMS week. New lasers presentation.


15 Feb

EN-EL intervention start.

Was missing fire detection schedule. and not informed that power cut will affect also the others laser rooms !! power back Friday 17. Switch on laser 1 friday afternoon for the weekend. see elog.


14 Feb

CV  intervention done: air conditionning, cable control..



8Fev New laser:


Final layout UO 213.pdf

laser1-2 OFF , magnet test,ect.


1 Fev

Laser1: calibration ok.  21A, 40% attenuation. Noise on laser23 histrogram display, strange.

Laser2: 24A, 70% attenuation.


31 Jan

Laser1: flow tube broken. new lamp, new gold reflector. ETM=- 5570hrs.. ??

Laser2:  no tuning.

run the night OK.

PC: kernel version problem, can't start ./laser23 ! kernel should be xxx89 cernsmp and not 103. Wrong Kernel by default now after booting !!!

no cooling before 16 January



16 Nov

Laser1: change internal attenuation from 20 to 40% (increase power) requested by Federico. Just for 1 day..


11 Nov

B= 3.8T ~20h evening. laser1 and 2 ON. laser2= 24A, laser1= 23A.


7 Nov

Laser2 main laser.

Installed second pin diode , the same as for the Matacq now (12V)



B field = 0 T 1Nov evening --> see difference of FWHM for the 2 lasers before and after !! see the matacq plots 1 and 2 Nov

Laser1 as main blue laser (maintenance laser2) 23A for Saclay..as requested. 60% intensity

Laser2 as red laser, second laser. 24A. New lamp.

Lasers OFF the 3 Nov : no cooling. Lasers ON 4 Nov for the weekend.


18 Oct

Laser1: change attenuation from 40 to 17% (ADC 500 to 250 count)

Scan of intensity and FWHM ~ 1 week


17 Oct

Laser1: error understand. Water connector YLF damage inside. no water flow on Qswitch and LBO thus heating (error and LBO mis-alignement)

(Laser1 stop Friday 14  afternoon while there was no trigger.. then again friday evening...)

+ New HV pulser  ! at the same time found TiS blue trouble, pulse not stable, ect.. : no more HV (cross IR..). Ship for reparation..


12 Oct

Laser1 back to normal, stable RMS intensity (tuning of HV pulser, decreased a bit power) current 21A is OK !


6 Oct

Magnet discharge 7.40am. lamps not OFF thus not a fast discharge ?

Laser1 new lamp, new crystal. error Qswitch fault.. see Reports laser1_11Oct2011 OM meeting


5 Oct

laser2: for the night change internal attenuator from 30 to 99%, and remote one from 10 to 3%. test for Matacq linearity..

Previously, day before, a scan has been performed changing only the remote attenuator .

JLouis: inserted asplitter for the Matacq cross-check (few plots out of spec)


26 Sept

laser2 small TiS tuning because RMS amplitude drift.

decrease attenuation after lunch (call from DoC):  25 to 30% to get green sequence on the Matacq. adjust delay

laser1: decrease current: 23.7 to 23.4A and adjust delay.


21-22 Sept

laser2: go to 26A ~ 1 day, then 27A. go back to 24.5A.


15 Sept

laser2 online instead of Laser3:

- 24,5 A

- changed configuration file laser_h.cfg

- changed the switch 1x5 mapping: channel.map2.  when laser3 is called, channel 2  is selected . No hardware modification like this.


11-12 Sept

Blue laser1 disappeared Sunday night> No investigation from on call shifters.

Monday morning: just put internal trigger , tiS box in manual , nothing was wrong. mystery. no error on log file.


9  Sept

friday: put feedback ON


31 august

Laser1 New lamp



- TiS shutter3 GPIB error, from 19h28  25 August up to 11h 26 August, time when ECAL call me .

DQM shifter didn't call the laser expert in this case can't react (shifter)



- stop at 16.24 the Neslab to see difference because yesterday no difference with Nesalb on the matacq.


22 August

- put spare Neslab , 30degree, laser1.  Neslab was OFF before by Adi (temp. unstable, 37, or 40 degrees)

- put air unit ON (after power cut was stay off, only spare air unit ON)

- delay adjustement:

laser1: 8,075 to 8.07 us

laser3: 7.94 to 7.935us.



neslab: found at 37 degrees following blue power degradation. As friday 22 July.

see e-log 639233

laser3: RMS amplitude above orange line..change current from 20 to 21 A.


22 July

holidays.. problem with laser1: drift amplitude and then saturday switch stop : problem TiS box shutter.

laser1: 23.7A then sometimes the week put back 23A, no more feedback (OK)

no laser during 18hrs. first e-log  620296 



Laser1: change internal attenuation from 50 to 35% . ~320 ADC count now instead of 450 (see plots matacq)



Laser1: new lamp. small tuning.calibration. (safe before holidays)

Laser2:  no blue light. mis-aligment always..from start up. No spare laser bight or weekend.



B field ON

laser2 no more blue. need tune the LBO.

laser1: 50% attenuation



saturday fast dump. Monday evening: laser data, no B field.

laser1: 23A. A=8.09us, B=3.2us. 60% attenuation not yet maintenance.

laser3: 20A, A=7.94us, B= 3.34us. 10% attenuation

received 2 YLF LBO, 1 valve, 1 gold reflector from Europe.


8 July:

make some RUN  with laser1/laser2 and with prepulse filter or not.  e-log: 599933


Tune TiS3 to improve power (first calibration too low)

Tune TiS2.


Laser1: just check main fiber coupling and then power OK on Matacq.

Laser3: new crystal !


4 July

tranfo to 207V

laser2: still not OK

problem: was the logic unit.

the 2 power supply are OK !!!

--> spare logic unit 234 : very low power for YLF2 while the threshold is ok.

wait for the 224 logic unit.


4-8 July

Plan for Liyuan visit:

- laser1: small issue jitter with external trigger mode

- laser2: tuning

- laser3: tuning. board 234 investigation if time.

- prepulse ; jump effect: try test to modify pulse raising time.

- all laser: check pulse on matacq.


29 June

B field = 0T

received repaired power supply unit. could test laser2 with no field


23  June

Laser1 no problem, online..



Second green filter mounted on laser1 output (the one used on laser2)


Monday 20 June 2am:

Laser3 failed: lamp and flow tube broken. Unscrew YLF chamber...

- 21 June:

see water leak. One screw on YLF Crystal not well fixed..

Tuning : power not very high, problem somewhere. maybe the crystal.. 620mW @ 20A

TiS red: can't get pulse stable. instability. FWHM too big also, 45-50 ns now ! Run at 21 A !! YlF too low maybe ?

15 or 16 June:

the new power supply laser2 failed again when tuning TiS ! Electric arc on main contactor inside the board.

measured phase and try to investigate with Sylverio. 235 V between 2 phases. According to Quantronix should be 190-215V.

BUT most probably: B field. the girder is just on the roof, the air-unit failed on this room also in the past..


10 June

Laser1 ONLINE : rms amplitude OK , problem disappeared.

On Matacq: low power, ADC count only 200-300 even for the RED while I didn't touched it (only new logic unit). need to go back to original remote attenuation which was 10% and not 15.


8 June

Laser1: not yet online because internal trigger all OK. external not --> double peak timing. New DG535 didn't help. not the problem.


7 June

Laser1: new HV pulser and change pockel cell also. See calibration. run ongoing.

Laser3: current jump from 20.5 to 22.8 by mistake. few hours. but ECAL saturate with RED , with this new logic unit.. need to change remote attenuator.

6 June

Laser2: online OK, start to drift (timing, power)
Cooling valve OK while it is logic unit from Laser3 which make valve trouble.

Laser3: valve OK thus not yet conclusion..

Laser1: change the lamp, tuning. small TiS tuning. seems OK.


2 June

Laser2 ONLINE because of laser1 degradation. small Q-switch tuning 1June, power increase but over night unstable...

Laser2: 25 A, 99% TiS attenuation, B=3.1us. Remote attenuation change from 10 to 20 then 30%. plot RED in Matacq but normal, power too low, but OK for ECAL (attenuator after the matacq)


1 June

Logic unit:

- the one on LASER2 : RS232 out of order, this is sure (not working on laser3 also). to be shipped to Quant.

- the one on LASER3 with valve problem: put on Laser2

- the spare one, model 234: put on Laser3 because no communication (to be investigate by Kejun on spare PC)

Check power / current with this new logic unit 234:

RED laser power quite different on Matacq:  put 20.5 A on board. control box with Key not working..


discussion OM meeting to see how to define laser HW intervention.


Laser2: see that the green filtre bandwidth is bad --> green light after the filter unlike laser1. replace with new one.

check power: no filter 32.7 mW, previous filter installed up to now: 30.7mW, new filter: 31mW.


25 May

Continue the work on LASER2 TiS up to now. still not OK (see histogram)

Laser2: 35 mW @ 23 A after tuning from scratch. T(IR) = 16.2- 16.5 A with no PC (no pockel cell)

Laser1 almost OK .

RED laser3: Valve ok from last 4 days.. even if there is some noise during operation.


17 May

LASER1 online, small tuning. to work on laser2.  attenuation: 35%.   No more RMS jump.

Laser2: 40m@23A, YLF: 1530mW@23A. YLF too low: was 1880mW before dismount the YLF chamber (flow tube broken)



Try LASER2 online, 11am. then move to laser1. too low power on laser2 (below the ADC limit)

LASER3: low YLF. need service. see on DSO before data taking.

Laser3: new valve problem: stay open all the time as the 5 of May. Power on wire is 15.6Vac, instead of 24 Vac (as laser1)

after switch on/off laser, leave the valve cooled, all is back to normal.


Laser1 too high power: move attenuation from 50 to 35%.


13 May

Laser RUN restart.

laser1 online, 23 A 50% attenuation. power: 450 ADC, RMS 3.5% but Seems stable . No huge jump at the beginning of the sequence unlike before. Tuning seems help.  FWHM ~28 ns (was 23)

laser3 online, 20A, 14$ attenuation.   power : 550 ADC


12 May

Laser2: small tuning again.

Laser1: TiS tuning . try new poco cell. no difference on DSO. get tuning to have pulse stable. need to see over time how it goes. On slow monitoring: can see noise on the TiS pulse.

Laser1 online


11 May

Laser2 done. power not so high, closed to the ADC limit on lightchecker but pulse very stable on DSO. FWHM ~28-30 ns but need to work on laser1.

Laser2 online. to work on Laser1 (huge RMS, sequence not so well)


10 May

Laser2 : tune YLF from scratch because pulse shape not so good. Found a piece of glass inside YLF chamber thus dismount everything (take out the crystal). change lamp again (was new but already visible damage !) Check and tune IR from scratch. Power lower than in March but pulse OK. (1600mW@23A now, was 1800 March).

all with B= 0T.

TiS2: remove Q-switch to re-align correctly (pulse not stable). all OK at the end but FWHM=29ns. Need to improve again.


9  May

Laser2 investigation: can't turn power ON with key. --> solution is to replace the full power supply 10K.

After that switch on laser is OK. start to process some YLF tuning (1600 mW@23A but was higher)

Tune TiS but FWHM >30ns so need to continue, pulse not good (was stop by a cavern evacation due to B ramp down)


7  May

Laser1 online by shifter because Laser2 failed.  Laser1 show again power RMS unstability. need to replace q-switch or HV pulser.


6  May

Laser3: need delay adjustment. internal attenuation from 17 to 14 %.

Laser2 OK

Laser1: tune TiS because internal run the night show power instability from Q-switch/HV


5 May

Laser3: power degradation over time. need maintenance: change the lamp after normal operating time (was waiting the technical stop but failed before !)


4  May

Tune laser1 while laser2 online for a while. YLF laser1 was mis-aligned after that power increase , OK and TiS very stable (2ns jitter) so put it online to see how it goes. 


3  May

Laser2 out of order. error. can't restart: change the lamp then OK ~ 15 min and suddenly lamp and flow tube brake. Huge water leak.

Laser1 online. rms amplitude not stable.


15 April

B field ON again.

Laser2 ONLINE  because Laser1 RMS instability. tuning didn't help (pulse ok on scope..) 22.5A, 30% attenuation.

Laser3: was OFF. turn on back. small tuning.


4 April

Laser2 OFF --> water leak--> pipe connector unscrew on YLF chamber. Try after: no more green light. Need tuning from scratch.              B=0T.

Laser1 online (still technical stop)

Laser3: small tuning, check coupling main fiber.


1 April

Laser2 Online during Technical stop , still ongoing: 22.5 A, 50% attenuation.                       B=0T.

Laser3: 20A , 17% attenuation instead of 10% ( low on Matacq, lamp 700hrs..)


28 March

Technical stop B=0T. Quick tuning laser 1-3. laser3 new cartridge.  lasers still very stable. 


18 March

Laser1: current was 23.6A with Feedback. Move back to 23A and re-adjust timing A to A=8.145 us (was 8.16us)

- requested by RFM but not necessary from Saclay : reduced intensity again because average ~3500-4000 ADC...

Now internal intensity reduced again from 30 to 25 %.

On MATACQ : power  ~ 350  ADC for blue is now reference.


15 March

Request from P.Gras to decrease blue intensity by 20% because of saturation : internal attenuation Laser1 from 40% to 30%. 

Light checker power was ~550-600ADC, now 450 ADC counts. Laser1 power was like this since 4 March so saturation since a while now.



11 March

During holidays, need reboot program , seems switch was blocked on the same position...



3 March

- Laser1 online is OK . no more waves - unstability on the amplitude RMS. see matacq plot.

- Test coupling for APD/PN ration with laser1 main output fibre. 5 x posistion.


2 March

- Laser2: low water level : leak, water on the YLF laser while the  lamp is ok, nothing strange. remove and install back the same lamp.. check the screw.

- Laser3: "current error".. --> lamp broken one more time , in the middle as the 4 Feb.

reason unknown but all was new (housing, ect..) . maybe the crystal is moving (unglue)  ? ..

Replace the TiS main shutter : not working in manual mode. not stable in Feb. time to time also (lost first event laser time to time)

Laser3 tuning, quick YLF and RED.

- Laser1 ONLINE


1 March

- Laser1: Change the lamp; YLF tuning; change the cartridge filter and the temperature sensor because of Di Water and Water flow error all the time.


22 Feb

- Laser1 --> Laser2 ONLINE because of instability on the energy RMS (other OK).

no data from 22 to 24 Feb (ecal OFF, test N2)


14 Feb

- Laser3: main shutter stop working --> no red sunday-monday morning

- received YLF spare part (20 lamp, Xcal, LBO, reflector, polarizer, spherical)


10-11 Feb

Laser1-3 online data taking with B field 3.8T.


8-9 Feb 2011

- Laser3: new YLF chamber because need to dismont everything : flow tube broken inside. found the crystal metallic end face is unglue or closed to be.. see calibration.

9Fev: check again YLF IR: IR power increase but not the green.

TiS : replace HV pulser --> why the reason why pulse not so stable (3-4 ns) 


26 Jan 2011

- calibration done.  Good health check with laser.

19 Jan 2011

- Tune and test  laser3.

18 Jan 2011

- Restart laser1. calibration.  internal trigger. (no B field)

12 Jan 2011

- End of cleaning all optical connector : switch and main laser fibre. Channels involved are here.



8 Dec

End of 2010 operation. stop everything IP5.


5 Dec

Sunday, B field stop so lamp switch off. change laser2 to laser1 (no remote control laser2..)

Monday change back laser1 to laser2 because laser1 not so good.

Laser3 was put back at 20A but was not the last value  (was more 22.5A) so delay change..and then no remote control to change the delay (instability RMS higher)



- Laser2 Online (RMS laser1 too high, outside spec.). increase attenuation: 35 --> 50%



- Laser1 setting: change 35% to 60% and 22.5A to 23A

- Laser3 setting: change 7% to 12%

Laser3 : no more data on Matacq while there is still good pulse on Slow Monitoring. Quick services (because run with ions..): very low power on RED main output (below 50%), YLF quite low but Ok for the lamp 1000hrs. Change the lamp then found TiS Q-Switch mis-aligned. Strange. Power recovered, OK. See plot.

Laser3: 20A, attenuation go back to 7% --> ~500ADC on Matacq, OK!



- Laser2: new lamp, usual maintenance.



- Laser1 online



-Laser3: No tuning of RED laser but RMS much better. (matacq plot on histogram)


1 Nov

Laser2 internal attenuation 40% to 60% to be far from lower limit.


spare part updated : received 6 Phase modulator for the power supply.



- technical stop ; No B field. then B field the weekend, stop again ; ramp up Tue.27.

Blue laser2 out of range. Vladlen change laser settings (timing, attenuation, current)

- Run Laser2 internal trigger 1hr 22.5A: ALL OK. power slow monitoring as before.



- Laser3: power decrease (ylf) . Change the lamp (500hrs..).



- Laser1 calibration

- prepare test with ecal-laser-room-05 to check GPOUT trigger of the optical switch



- Stop laser1. Time to change the lamp  1000hrs See matacq PLOT.

Change with Laser2 online  

- Problem restarting laser2 and also laser1 after new lamp: 3 phases monitor died, the 2 lasers at the same time !!

- Problem with VME interface laser supervisor. Clean fiber of PCI board.


27 Sept

- switching OK with switching time from 8 to 4sec; log file now from slow monitoring (used for light checker)

- today switch from ./laser22 to ./laser23 removing the intermediate position 0 of the 5x1 switch. Now optimized when moving from blue to red.

- LASER3: best jitter - tuning since a long time after dismount  YLF chamber (flow tube)


23 Sept

- Laser2: quickly used it online - test ok data OK

- Laser1: current were 23.8A- change to 23A and re-adjust delay A. Laser still stable, ok.


22 Sept

- no-handshake mode OK

- switching time reduce from 8 to 4 sec ok: sequence is now 38 min instead of 48 min.

since RUN 146325, e-log ID 449159.


21 Sept

-Laser supervisor:  removed handshake morning (acknowledge switch position) and afternoon reduce switch time allocation from 8 to 4sec (average time on laser22 is 3.5sec)

now,  error in case of switch or shutter problem will be visible on Mataq plots only. no more on laser supervisor log. The switching time is based on laser log file from ecal-laser-room-02 (copy every few minutes)


20 Sept

- Laser3: YLF died Saturday (plot dso21664-3.ps). Found flow tube and lamp broken (600hrs). Lamp brake because of flow tube; Put laser3 online immediately after. Change internal attenuation from 10 to 7% because of + power (RED 143mW@20A no attenuation)


31 August

-laser1: replace flow tube, because closed to be broken (crack). new lamp (1000hrs)

Online for ecal without B field to test Gain 200 (change 0dB to -10dB)


30 August

TC, technical stop. Magnet ramp down. Maintenance/checking  of laser1-2-3. laser1: instability last hours.


25 August

- laser3 back to calibration sequence. Issue with RED laser: long switching time from blue to red or EB to EE so lose 1-2 LM region. Also after some times no respond from Laser22, no more blue and RED. Restart remotely laser22 OK but freeze after 1-2 hrs... Found underground ICS box laser3 no more display and DG535-2 (spare) wrong trigger rate. RED laser was lost after B field stop and start so could be the power supply of the ICS box as already suspected.


23 August

- RED laser: change the lamp because impossible to restart laser after B field interlock (mid August). Laser back to normal despite lamp not broken.

- laser1: no problem during the last 3 weeks.


21 July

Refresh training in the laser barrack , with Chris Jan and Vladlen.

Turn on laser2 to check laser.

Updated TWiki page.

No B field from Monday 19 to 26 July.


19 July

CMS techincal stop so change lamp of laser 1 and laser 3 before holidays. No tuning of TiS.

- laser1: start at 22 A, 35% attenuation as before

- laser3: 20 A, 10% attenuation as before.4

Replace ICS box of laser3 because one more channel RS232 out of order (only channel 97 is working !)


18  June

LASER3: "current control error" --> water inside the lamp !  Lamp 400 hrs...

Full tuning of YLF and TiS.


14  June

LASER2 out of order after weekend (current = 7A, not in used) --> flow tube broken, lamp ok, no leak upfully.

Switch to Laser1 ONLINE because needed to remove flow tube, unscrew YLF chamber, repaired soldering connector power supply lamp...

Laser1: 22.5A, 8.39us, 2.9us, 35% tis attenuation

Laser3: 20A, 8.13us, 3.15us, 10% tis attenuation


3 June

LASER2 and LASER3 start online: adjusting setting:

- laser2: 22.5A, 8.22us, 3,10us, 40%

- laser3: 20A, 8,15us, 3,15us, 10%


2 June

Laser2 and 3 ok after night run internal trigger. Wait B field ramping again after CMS-LHC stop.


1 June

change Parker Valve again with new one (yesterady replaced with old valve already in used )

and OK. so valve is not working , dust.

Clean water tank , replaced distilled water.

Run the night 1-2 june laser 2 and laser 3 internal trigger. Plot on histogram. Lasers OK.



Pt5 power cut and also cooling during the weekend; the 2 blue laser OK (but not restarted during weekend because of wrong ./laser22 selected)

- laser 3 : lamp broken; and strange behavior of secondary cooling water: try new temp sensor, new Parker valve: no effect. Flow meter shows always open circuit, like valve open on laser cooler. But valve on/off seems OK (noise) and was of few minutes at laser on.



- Laser1 stopped: Water leak --> Flow tube unglue , out of the YLF chamber !  lamp broken...

Switch to Laser2, online : 22.5 A, A= 8,39us, B=2,9us.



- Laser3 : YLF and TIS fluctuation (timing); re-adjust timing remotely oevr the weekend  but YLF recover iteself after few hours; last delay A = 8.17 us, 20.5A

Increase TiS attenation : 8 % --> 25%

- Laser1 still OK, feedback on (double peak on TiS timing but OK on Matacq, too small effect)


4 May

- Laser3: change TiS box. because Saturday, lost control. (green LED not OK time to time)

- Quick tuning laser1 and laser3 :

laser 3 jitter not stable; tune HV and seems better (pulse shift but ~50 ns, stable HV )

Laser1: fwhm goes from 23 to 28 ns, jitter ~8ns.



20 April

- Matacq Laser1 : Average RMS amplitude and jitter OK but spread. Slow monitor: 2 peaks.

- got 2 ICS box controller as spare;

- Attenuation , power scan (Philippe)


19 April

- quick laser1 tuning, because RMS amplitude reach 5 % (others are fine) tune HV pulser, was not tuned when I changed it.  23A --> power = 103mW (was 88 13 April)

Delay A=8.38 B = 2.9, 35% attenuation (was 40) for 22.5 A

- New ./laser22 code : add information on the shutter , to get real status of the switch position and not theorical one.


16 April

- Laser3 online


14 April

- Laser2: problem of last weekend was the lamp --> broken , on side brake when I pick up from YLF chamber, water inside.

Lamp: new model from Quantronix, only 439 hrs !!!!


13 April

- Laser1: new HV pulser. Was faulty (check with polarizer), New calibration reference. Working curent = 22.5 A.

- Laser1 online at 17h20 : working current 22.5 A, delay A = 8.39ns, delay B = 2.90ns.

Can start laser2 investigation. 


12 April

- From Friday evening: scan with remote attenuation, P.Gras, all the weekend.

- Laser1 with internal trigger stopped after 8hrs (friday evening)

- Laser2 : YLF change a lot ! why ? no reason. Was very stable before, I touch anything. Trigger ?bad control from Philippe on GUI ? trigger?


9 April

- Laser1 calibration done with new crystal.

- Laser3: check tuning. all ok.

- received YLF LBO from reparation (25 Jan, laser1)


8 April

- Laser3 back Online; was removed since 25 March following calibration problem and air-unit installation.

- Laser 2 still online, feedback : power at 22.5 A now.


7 April

- Laser1:  New Crystal Rod --> found 3 spots on crystal (was installed since 2007, before USC installation)

Laser1 was still unstable so last thing to be checked was the crystal.


26 Mars :           

- Spare conditionning unit installed; RED laser out of the run for that and to understand problem: switch moving while no trigger !

- New lamp Laser2 and Laser3. Laser2: working current 22A, 45% TiS attenuation (run dso21446-2.ps)

- no more GPIB error.

- Laser1 YLF tuning again, plot dso21448-1.ps


10 Mars :           

- long run 4 to 10 Mars: ./laser22 not answering 10Mars midnight. No more control, gpib error TiS3. Memory problem? run too big ? (200Mb .dat file)


3 Mars :           

- GPIB error but for TiS laser3. Try spare TiS box but communication failed; same on laser1 --> was the repaired tis box received in Dec 09 !


2 Mars :           

- GPIB crashed, after 4-5 days ; error: TiS 2 ERROR  investigate hardware and found connection to shutter not well connected on TiS box --> re-soldering. Was the problem (operate shutter manually I found the problem, the old shutter I've replaced was thus OK)

- Go back to 2 PCI board and new software, go back with AC smart unit control.

- tune laser3 in the meantime (jitter was 6ns..)


1 Mars :           

- switch_test# ./laser22 still in live; no GPIB error ! run more than 3 days OK ! (dso20429-2.ps) 

Thus problem was ICS box with AC smart power unit

- Can go back to 2 PCI board for next interruption (maintenance)



- GPIB crashed again so problem is not 2 PCI board.

- change the power supply of the ICS box controller : removed power from the AC smart unit, put back usual power supply with new Adaptor 24 V (was 12V before)



- feedback laser3: very BAD result for jitter--> wrong setting for laser3 (was blue setting instead of red) see Matacq plots.

- Students interventions reports here



- Problem GPIB every night; since 8 Feb if there is long global run... change ICS box, cable...didn't help.

--> Go back to previous software version with 1 PCI board only (keeping new PC) --> #switch_test

- Put feedback ON for laser2 and laser3.



- Laser3 fine now after YLF tuning --> no more huge RMS. Matacq OK, as laser2.

- GPIB crashed again 2am but error on TiS 3 , was on TiS 2 before swapped the ICS box so seems the ICS is faulty.

Replace with new one, spare (need order another spare one)



B=3.8 T, start run (not yet LHC)

Swapped ICS GPIB controller because laser2 GPIB TIS failed (long run)

Tune YLF laser3 (rotate Qswitch), calibration online.


9 Fev:           

GPIB error, crashed the RUN again : 3rd time for 3 long run !!  (more than 1 night)

Crashed last week and week before during the night (Kyoungha) Suspected NEW DSO.


8 Feb :          

Start Global RUN for indefinite time, Laser2 Laser3 online


28 Jan :          

Laser3 stopped during global RUN (2 days) --> faulty Qswitch (RF driver error as 15 Jan, new RF driver didn't help)        internal attenuation = 8 %

• Send Power supply and Cooler assy. from laser1 to Quantronix


26 Jan :       

Change PC ecal-laser-room-02 with new one.


25 Jan :       

 - Laser1: back to normal after following intervention:

                    • new lamp, 24941hrs

                    • replace YLF LBO because of water connector

                    • New power supply and new Pump cooler ASSY because of burning connector (transformer)

                    • Try new RF and new Qswitch but YLF pulse not stable on scope. But TiS back to stable after 1 week....(poco Cells tuning)


15 Jan :       

Laser3: faulty RF driver, replace with new one. BUT after tuning screw on RF is back to normal so re-installed original one.

14 Jan :       

- Laser2 online local run (CMS open) 23A (should be 22.5) internal 25%, A= 8.41 us, B=2.9 us, timing = 11750ns.     One shutter OUT (main TiS shutter). replace with new one.

- Received spare parts and new IR detector.

5-25 Jan:       

 Work on Laser1 because last year YLF not stable YLF was 12 ns:





14 Dec :        - Laser1 : RMS jump : from 4-5 to 7-8 . Try to tune laser 1 but problem is YLF RMS (~17-20 ns !)

                    Need to tune it from scratch --> wait next year no B field to work with IR viewer.

                    Put Laser2 online 22.5 A internal 25%, A= 8.41 us, B=2.9 us, timing = 11750ns

                          Laser3 online 20 A    internal 10%, A= 8.04 us, B=3.2 us, timing = 11750ns

                    - Fine tune Laser3 (power went down)

                    Maybe come from air unit on control room which was stop (but since friday)  was 24.5 degrees.


8 Dec :          First try with LASER3 online: 20A, internal 10%, A= 8.025us, B=3.2.us, timing = 11770ns

                     Laser1: attenation tis = 30%, timing = 11780 ns


4 Dec :          Laser1 power too high: internal attenuation reduced from 50 to 18 % ! on Fast monitoring is ~4-5 ADC  (was 7-8 for laser2) --> MATACQ is 300 ADC as laser2.


3 Dec :           break ecal run so: put Laser 1 ONLINE (22.5 A)  and Laser3

                     Need to check timing, ect. after Liyuan visit (timing configuration, ect.)

                     - quick tuning of Laser2 --> get back good jitter and good FWHM (<25 ns)


                     on Acqiris: # ./monitor 11700 should show a pulse at ~150 based time for all lasers (ext. trigger, gate D no more in used)


                     - intalled AC smart Unit


20 Nov :         Friday evening: first LHC beam. Magnet CMS OFF


13 Nov :          Laser1 error (YLF leak) Switch to Laser2 ONLINE 22A, 38% internal attenuation, timing= 11765ns

                      Laser1: new YLF chamber (was still the original one, 1999), new flow tube, new lamp.


10-20 Nov :    Liyuan, Kejun: new software, 2 GPIB card,ect. No tuning at all on lasers.


2 Nov :           LASER1 lamp brake Sunday (B field stable at 3.8T) time = 1770hrs.

                      Calibration: dso21305-1.ps

                      Setting: 22A to have FWHM= 30ns, 50%, A= 7.42us. Fast Acqiris = 7-8 ADC (~13 ADC for 100%)


28 Oct:           LASER1 jump --> YLF jump, due to B field ? run DSO21280-1.ps stop 28Oct at 14h30.

                       Should be no trigger - no ECAL during B field increase ??? look timing of laser plots..



27 Oct:           B field ON, 3.8T


25 Oct:           Laser1: check and put back 23A because claim power is too low. A=7.42ns; power = 0.33mJ.

                       check attenuation: 13 dB (20m long fibre). OK. Quick tuning, RUN : dso21279-1.ps


22 Oct:           Laser1: global laser not good so previous tuning was bad. New tuning very quickly :

                       current = 22.5 A to have FWHM=30ns, 50% attenuation OK (30mJ histogram), A=7.40us. DSO21264-1.ps


19 Oct:           Check laser1 and laser3, 30mW@23A for the operating current: calibrate power on slow monitoring (30mW)

                      Laser1 with blue filter: no difference on power so leave internalt attenuation laser1 50%, laser3: 20%


13 Oct:           No laser all the week --> cms off, water leak.

                      Clean/polish main fibres lasers / optical switch 1x3. (3 lasers) all were not perfected and need small polishing.



12 Oct:           Laser1: put blue filter, FD1B as for laser2. Move main SHUTTER because even if it's closed pulse was not in the middle and there was some diffuse light after !


2 Oct:             Laser3: internal attenuation at 20%@800nm to have ~30mW as laser1.

                       Change setting online for LASER3: attenuation = 10% and delay +20ns so 11770ns (as seen by DQM). don't know why for the timing. On Matacq, difference of amplitude comes from pin diode DET10A (x3 for RED / blue)



1 Oct:             Laser1: internal attenuation at 50%. Remote attenuation has been changed on DQM.


22 Sept:         Laser1 current scan to get settings delay and attenuator . Tune Laser2 for GREEN (calibration)


17 Sept:         cooling back. Restart LASER1 online  (small tuning of YLF and TiS)


7-17 Sept:      no cooling, no pc: pt5 maintenance.


30 August:     LASER1 lamp went OFF as foressen with B filed interlock  (slow dump) - sunday evening.


21 August:     Laser3: new lamp, calibration, tune YLF and 800nm to be ready for ecal.


20 August:     Laser2 calibration: change all YLF chamber components. Working point: 22A, FWHM=23.5ns.

                     Power: 30% more than last run online at 23A. A=2.9 us.


18 August:     LASER1 online. Because Laser2 lamp broken - same time as magnet ramp down (~7am). Flow tube broken also.


3 August:      Summary from 28July (not here and B field: start CRAFT)


31 July:        Laser2 problem the week: 1) GPIB crashed , new DSO... 2) air conditionning stopped Pulse drift....

                    Feedback ON why ?... 3) Pb with DG535-2: need to be reboot and error communication 4) Lamp OFF due to bad info on magnet interlock.


16 July:        TiS box LASER3 repaired --> soldering of green LED.

                     will changed the trigger for Fast monitoring --> trigger selection fine with internal mode but not External !!

                     - Laser2 OK on ECAL: ~2000 ADC for EE (0dB and 50%) and ~6000-7000 ADC for EB.


15 July:        Check LASER2 power --> lower than 7July on EE according to S. Ledoskoy.

                    - Full chain: found again dirty fiber on optical switch 1x5 (polished again)

                    - change attenuation box with old one from H4: positive pulse is DET10A and no more DET200 (obsolete)

                    - TiS2 power on laser: 44mW. On Ch99: ~3 mW so 11dB. Just before intervention was 0 mW !!

                    - Fast monitoring: 8.5 ADC count for laser2 with green filter, 10-11 ADC count for laser1 (no filter), 36 ADC  for 800nm


14 July:        LASER3 ready to be ONLINE. Online with new digital scope. Should not crashed anymore !

                    Keep LASER2 for ECAL , with green filter.


9 July:         - Magnet field ON, start. 3 Lasers internal mode. 

                    -> LASER3 to be ready next week.


6-9 July:  Tune Laser1 and Laser2, Delay A. Laser1: remove p.cells to check power. Continue polish fibres on laser : done for slow monitoring, laser 1 and laser2 with new polisher.


24-25 June: Tune Laser1 (YLF also). Got new polisher.


18 June:      LASER2 online with green filter to tune LASER1 (tune on friday 19)


17 June:      LASER1 online with green filtre from 4pm. But put shielding box on YLF before --> YLF not stable, jitter=7 ns.

                   Matacq plot:  Matacq_18June_09.pdf


9 June:       LASER2 online to put green filtre on LASER1 --> removing pre-post pulse (post pulse see after 4us)

                   Check tuning LASER1 --> it was bad because lasing without trigger ! --> explain high power !

                   Tune TiS to have no lasing up to 23.7A so OK --> more stable (dso20910-1.ps), jitter=4.5ns @22A.


8 June:       LASER: internal mode after small tuning to check stability: dso20909-1.ps, 22A


4 June:       LASER1 : power went down mid night. Same time as current 22.1 --> 22.2 A. YLF increase but power TiS decrease !

                   Files is DSO20890-1.ps. --> temperature of the room ?... Tune again Qswitch laser1 . Power is back.


2 June:       LASER1 online for Cruzet but begining of run -> wrong settings. all delay = 0 ns, TiS attenuation=40%.

                   After a while, power very low, and 3 JUNE; no more power (very very low)

                   --> Tune TiS laser again !!! (Qswitch)


27 Mai:       LASER3 TiS box: 1 pin RS232 not connected properly inside: wrong jumper in used.

                  Adi: monitoring study : found signal very low after laser trigger, after ~4us after TiS pulse !


26 Mai:       LASER3: tune. 115mW@20A. TiS box not answering to GPIB. To be investigated. Used a spare one.


                   Tune Ti:S Laser2. best FWHM=27ns @ 23A, power=60mW. Max power = 70mW with LBO alignment

                   but FWHM>30ns. (up to now FWHM=32-33ns)


25 Mai:       Put LASER 1 online --> P. Jarry, to see difference fwhm,ect.

                  Tune Ti:S Laser1. Working point 22A, delay A= 7.445us (fwhm=24ns, 66mW)


19 Mai:       Amplitude EB: ~4000 - 5000 ADC count , on fastcheck.

                  Before was ~2000 ADC, and last days go down to 1200 ADC once.

                  --> Saturated from P.Jarry (gain is switching to 12, bad) so need to decrease power by 30%.

                  Can't wait configuration file to be changed so changed internal attenuation TiS from 99 to 70%.

                  P. Jarry: nominal count should be 2500 ADC count.


18 Mai:       Check full chain attenuation with Laser2 at 23A: (because of P.Jarry comments, laser fluctuation)

                  1) before touching something:

                  TiS output = 60 mW; Fast Acqiris= 6-7 amplitude ; Ch.99 with 20m fibre: light but no power

                  2) Removing 1x5 optical switch --> clean and polish all fibres. Not yet perfect but better.

                  3) Found main fibre on laser side burnt --> polish from scratch by hand (polisher not working)

                  4) Power on Ch99: 7mW --> Lost = 9.3 dB,  OK. Fast Acqiris is now >12 amplitude (+40%)

                  Previous value measure:

                  2004: 10.5 dB; 2005: 8.8 dB; 2006: 10.4dB; 2007: 10.5dB; 2008: 11.1dB;


15 Mai:       Laser2 attenuation: get 6.8dB attenuation on first device, optical switch 1x5. Need to clean and continue next week.

                  Power TiS output= 50mW, power after coupling 1m fibre=38mW, power after 1x5 switch= 8mW.


14 Mai:       13-14 Mai: MWGR.

                   - LASER2: check attenuation, coupling of main fibre.

                  Can't see power value after optical switch ch99 but on Fast Acqiris value of the pulse is 7-8 count (OK)

                  Need to check full chain to recover power.


                 - DSO: put back old DSO online. Put Digital on with ecal-laser-room-01 for Kejun.

                  --> else all the time GPIB error during MWGR.


5-6 Mai:    ./laser22 crashed. log_5Mai.log --> HP2 head error then GPIB error --> no more control the 6 Mai 14h26.


29 Avril:    . LASER 1 : lamp broken after few hours only. Same as for laser2 6 March !!!

                  --> run during global RUN: Laser3 selected and  switch scanning but laser not needed !

                  . LASER2:  move mirror between YLF/TIS to allow space for box for shielding --> tune TIS from scratch.


23 Avril:     New cooling error. Laser2 stop. Switch to laser1 for 23-24 Avril. (Laser2 OK)


21 Avril:     Test new cooling OK. Leak detection OK. First long run with LASER2.


16 April:     Got power meter, laptop. Picture of IR CAMERA ok ( picture on calibration page for laser2)


25-27March: Laser2 Online, MWGR.


6March:     Laser2 error: lamp broken (250hrs). Also found 2 fuses broken on power supply and the orange phase monitor device

                  inside very noisy --> replace with laser3 one. Maybe lamp brake due to this problem.


                 setting: delay A= 7.55us and current 22A OK.


4-6 March: Global RUN: LASER2 online


29 Jan:      Laser request. Calibration LASER2 Friday 30

January    No laser (no cooling, no ECAL)

                 Cooling: wait new pump

                 Shielding:  investigate power meter and IR viewer.  Start shielding laser itself (A.Ball)






- From 21Nov: switch OFF laser up to cooling is back !

- 20Nov: pt5 meeting.  Waiting for action: shielding and DCS,....

- 19Nov: main power cut but transformer not affected ! why ? laser still on , pc 01 also...

Laser1 online

- 14 Nov: ramp up to 4T. measure current on cathode laser2, shield inside laser room.

. ramp up didn't break the lamp.

- 12 Nov:  shielding power meter MU metal didn't help


- 10Nov: LASER2 online.

- Saturday 8: LASER1 lamp broken --> magnet go down 0T. restart Laser1 and Laser2 Monday 10


- 6  Nov: magnet go down fast the night 5-6 Nov --> lamp LASER2 brake !!! as laser 1 the 24 Oct !!!

Field effect ??? --> on elog: Magnet fast dump the night 1h50 6Nov.

 LASER1 online because LASER2 lamp broken.


- Weekend 2 Nov: GPIB crate failure. communication. dso20529 (long but no history)

- Monday 3 Nov:  restart laser2 because GUI was OFF !

- 4 Nov: found that Feedback slope was 0 --> NO feedback ! put back feedback --> run dso20532-2.ps current >24A



- 24 Oct: LASER1: broken from 21 Oct ; flow tube damaged.  Green T = 7.2A but wait no field to be able to have power meter.

- 22Oct: today no field so tune laser2 OK.  Put LASER2 online because LASER1 failed the 21Oct --> lamp broken. Liyuan try to put laser2 online but failed.

Field effect ??? --> on elog: Magnet fast dump the 21Oct.

- 20Oct: discoverd with Liyuan than we can not used power meter and IR viewer under magnetic field --> can't tune laser properly !!!!

- 17 Oct: no laser events from 6am. Go the morning --> no more air conditionning Out of order.

Put LASER1 for the weekend  (no time to tune LASER2 --> run) Monday Liyuan is here

- 2 Oct: LASER2 stop --> cooling ion USC. At the start up, flow tube brake --> huge leak.

Change lamp and flow tube

Put LASER1 ONLINE but the Friday 3Oct LASER STOP and NO ACCESS to USC (but NO shift in ECAL for weekend)



- 10 Sept: found attenuation TiS setting at 40% instead of 99% (run dso20405-2.ps)

- 5 Sept: cooling interruption USC Friday evening--> restart laser

- cooling circuit done. Waiting for electrical power.



- Start closed loop cooling circuit on chilled water

- Continue access card reader installation (Cegelec)

- LASER1 : tune to be ready as spare. Center = 11750ns as ref.  (like laser2) Plot internal run online.

Change one connector pipe secondary water (leak, closed to be broken)




changed the 2 relay inside logic Unit (valve no more working) after cooling break at USC55 the week.



Laser2 leak again --> leak on YLF chamber itself. Changed with new one. tune YLF and TiS.



Laser2 ONLINE I=22A. dso20177-2.ps


25 April

Laser2: leak on YLF --> flow tube un-glue. New flow tube, new lamp. No time to run internal mode. Quick tuning.



1) Still problem with cooling.

Main interruption of mixed water.

When changing filter on laser2, laser1 body filter failed.

cooling pressure increase after restart !

Decided to installed on laser2 a stell body 30um filter (parker)

2)  Seems to be possible to used chilled water so used circuit from H4.


3-7 March Liyuan

Move laser1 at pt5.

New YLF Qswitch installed : 100mW so new RF driver installed (100mW) else pulse not stable  (old one was 20mW)



- Laser1: change HV pulser. Problem  solved. Laser1 back to normal. Shipped HV pulser for reparation.

"Water flow error" disappeared by increase realy switch from power unit 234(close to reset button)

- Same as for Laser3, I got same trouble.

- One scope failed in H4 during laser1 tuning.



Laser2: leak of YLF LBO: broken water connector (shipped to US) Tune all laser during all the week.

Need laser for EMTC, comissionning with few SM.

Test the timing of switching sequence.



Laser1: restart for Dee4 after cooling breakdown.


1) no TiS pulse with HV . change HV pulser, still no light Change Q-switch, still no light

OK to run 3-4 days without HV and delay, pulse still OK.

2) Water flow error time to time. Change sensor inside water tank, same problem.  but during weekend OK.





- 17 Dec: Laser1: Stop. End of 2007 EE laser test

  Pt5: Water more and more dirty --> degradation of water quality visible after 4-5 days of commissioning interruption.

- first test of Cyclone filtre not relevant --> filter very dirty after cyclone filter (picture)


19 Nov

Laser1: new lamp (2100 hrs). Quick calibration.



ECAL cabling at pt5. SM comissioning with laser2. Used EMTC (Mataq output for trigger)

- 2 weeks problems because of no air-conditionning (no TiS pulse)

- Problem with dirty tape water: installation of tank. Waiting cyclone filter

- 2 PCs have been moved to cms server (for communication)




by changing the lamp, found 2 spots on crystal surface and flow tube broken (--> wrong trigger rate ???)

Change crystal, lamp, flow tube, RF driver (because pulse jump, after 21.5A) and relay logic unit (water flow error)

Run at 23A: dso20096dat, FWHM=20ns.

ONLINE: start at 22A for Dee4.



Problem with external trigger H4:

complained that laser not working (during vacations) but found external trigger at 300KHz.

YLF was lasing but very bright green light and NO blue !!!



start EMTC installation (P.G.)



LASER1: new lamp. Max YLF=23.7A. After, no more pulse stable.



Re-installed Q-switch on Laser3: Q-switch makes same error as Laser1 "RF driver mismatch"

Installed new Q-switch --> OK

Received new RF driver, new power supply.



Laser1 online: stop run (dso20034) and feedback because increase to fast.

stop at 23.4A and go back to 22.8A.



received YLF spares:

- ROD Crystal. Data on calibration pages 1001-0955

- LBO, Q-switch, polarizer + fuses




LASER1: ERROR display (LOW RF) 391 RF driver fault.

Try with Laser3 Q-switch: same error.

Change unit with Laser3.  to be repaired. Tune YLF (check Q-switch)


5 to 16March

Liyuan and Kejun at CERN . Move lasers 1 and 3 at USC55.

1) Laser2:

- low power. find flow tube broken. Replace it but find crystal moving on one side.

crystal not glue anymore on one side.

Deside to move laser2 at USC55 and put Laser1 online for Cosmic.

- new crystal, golf reflector, lamp and flow tube

2) Laser3:


3) Laser1: online

4) New fast Aqiris: timing is OK now because A= T+ t(ns) for all YLF and B=A+t(ns) for TiS



Laser2: very small leakage


January 2007

Laser2 running back for cosmic calibration





14Dec 06

Laser3: YLF and TIS back to ''normal''

No WATER from 21Dec to 16Jan.


8 Dec

Laser3: water leak again on YLF chamber (as usual)

- New Gold Reflector

- New LBO assembly because of damage pipe connector (corrosion). Use spare one.

- YLF quite OK. Tune TiS.


4 Dec

Laser2: new flow tube -> broken inside but no leakage

Start Laser2 and 3 for SM recovery.


28 Nov

Laser3: replace broken flow tube. Life time: change after 1year and 3 month.



received new HV pulser as spare; new TiS box



End of TestBeam periodII in H4. H2 still working with laser (SM with HCAL)



New lamp: laser2&3.



Main POWER CUT at CERN. Restart Sunday and Monday (Checking&tuning laser2).



- Test beam started.

- Still Laser2 and 3 online.

HV from Quantum: no damage part. OK



- Laser3: change HV pulser with the repaired one (the original one,from 15January).

Should be removed Center instability (+-20ns)

Can reach FWHM 21 ns but some double peak at 26ns. Tune at 27ns to be stable.



- Laser2: new lamp

- Laser3: tuning because of too much current increase, feedback.



changed back /test4/./laser2 with the 2 lasers fired at the same time because of some missing RED events



- Laser3: YLF LBO water leak. This may cause from ~3-4weeks the pulse drift.

Lost pulse on XDAQ --> no red event.



- something no acknowledge from blue or red. Not stable problem

- check discriminator: ok (-800mV) up to 45% and 1% TiS attenuation.

Bellow: no more gate.


29May to 2 June

Liyuan's visit.

- 1x5 optical switch

- new attenuation box : one more remote control filter.

- new software: separate processing for hardware and data acquisition. 3 lasers online ok.

- tune lase3 (new lamp)

- spectral filters to check pre-pulses on SM.


23 May

power cut

New SM (SM4 --> SM12)



power cut - no water



- YLF3 interruption: LBO water leak (water tank level warning).



- new software feedback: ./laser2fb



- SM20 --> SM0

- change lamp laser2 (mistake because old one : ~350hrs only).

- Laser2 TiS: 60% . Power=37mW (Red=70mW)

- TiS center feedback for laser2&3: slope=-0.05A/ns

Test internal trigger: dso2932.



new history plots



Laser1: new lamp



- Laser2 and 3 online for SM20

- Laser3: new lamp, new socket because of corrosion. See picture.

- Check that energy is different on fast monitoring.

Laser2 TiS=99%

Laser3 TiS=40%



- Received from Quantronix TiS box3. Installed it for Laser2. OK

- Laser2: calibration done. Run internal mode. Working at 22A OK --> FWHM=25ns.



- Laser1 stopped all morning: water pipe connector broken (the one in plastic on power supply, input for YLF)

Buy and install a new one in stell.

--> LASER1 : Feedback on TiS CENTER



- new plots. summary from January.


8 March:

- SM19 online;

- Neslab : water consumption too high: cover need to be keep in place, solve.

Create instability twice: begining of March and 10 March.

- Feedback on TiS FWHM for all March. (jump to 22A up to  23A)


28 Feb

- new software, ./laser2test : H4 DAQ communication disable during data analysis.



- SM17 in H4 ;

- Laser 1 and 3 : new lamp , calibration.



- External laser trigger: 250Hz (seen by pcalth6. Create lot of file from Sunday)

Laser1: no more water in Neslab after 4-5days !! because of Hz ? (YLF: no flashing regularly)

--> stop lasers monday morning.



- first data taking SM16; safety seems OK for RED and blue even at high attenuation (2%)



- Laser1 &3 online (no yet received TiS box3 so used laser1 to use Tis box2)

ready for next SM: SM16 in H4.


15 Jan06

- received repaired HV pulser (laser3): spare part now.

- received repaired logic unit (cable broken -->connection): installed on laser1. OK

- received new power detector, new scope.


9 Jan.2006

- Laser2: new lamp, Ti:S tuning (removed PCell)

- Laser1 & 2: calibration

- send DSO to Liyuan.






15-16 Dec.05

- Send HV pulser 3rd lasers to Quantum

- Send TiS control box 3rd laser to Quantronix (Red led, make a spare one)



- Chilled water shutdown--> stop lasers.



- temperature laser2 room not stable from one week.

- Laser3 re-tuned: only HV unit energy goes from 20 to 80mW !

- laser2: little tuning (room temperature) 



- tunning laser2-3. laser3: HV tunning recovered energy from 56 to 80mw !


17 Nov

- Laser3: new lamp.

- Laser2 tuning; current 23.5A for SM5.



- calibration laser1 New lamp - new logic unit.


9 Nov

- received new logic unit with J105 cable. Works fine on Laser1.

- safety investigation (level2 TTL): level2 TTL not working properly. Take signal from the first cable of the MEM.

Problem of laser power: if too low, threshold hard to get (discriminator). For RED it's OK


7 Nov.

- Laser2 &3 running back ;

- Safety from MEM repaired:

1 BNC broken, 1 device coincidence unit broken, level2 TTL from MEM always signal --> use pin diode directly.


18 Oct.

- Laser1: no more communication (GPIB)

- communication back after power off all - check signal on RS232 and disconnect control unit.

- Laser2: control unit broken --> no more YLF control. Change control unit with Laser1 to have Laser2 online.

Laser2 ONLINE (need to change lamp of Laser1 after repairing the control unit)

- Laser3: some dust of gold in the water tank.


6 Oct.

- Red laser not so stable from now: stopped ecal DAQ to check lasers

--> 800nm: NRG=38.5mW@20A After tuning (HV p.cell) 79mW: OK.

To keep same energy level: TiS attenuation 40% to 20%.

HV not stable with time ? to be checked.

--> laser1: small tuning; all OK. 


- Feedback results on laser2 (run ~20hrs; count=200):

DSO2784: YLF energy

DSO2785: YLF center

DSO2786: Ti:S energy


26 Sept

- Red laser online

- Liyuan at CERN: investigate Laser2 YLF / feedback software on Laser2.


16 Sept.

Laser monitoring workshop. PDF on reports page.


13-14 Sept.

- Laser1 stopped because of chilled water maintenance. Change lamp. Make calibration.


24 August

- Laser2: change mirror HT/HR and LBO: no improvment, even worse.



- Laser 3: replace YLF chamber because of leakage.

- Laser2 maintenance: change lamp because power 40% lower (1000hrs) but with new lamp still too low.

YLF: 1400mW instead of 1800mW @23A.

--> Laser1 online for Cosmic rays.


1 August

Laser2 for Cosmic SM5:

20dB, 80% attenuation (99% TiS box), A=3.12ns

Laser3: not yet used because of flow tube broken.


21July 2005

- Lasers ready to be used for cosmic rays.


18 July 2005

- 1x80 optical switch arrived at CERN after mother board reparation. Now it's fine

- Changing of water pipe connectors (1/2" pt by 3/4" ht) for outlet Laser1 and inlet Laser2

- Water inside room1 due to big storm. Action to be done for the building


12 July 2005

- Laser1: stabilization of the YLF. Run of 20hrs fine.

- Laser2: outlet cooling connector broken (small leakage observed--> washer too old)


3 July 2005

- replaced all YLF lamp + calibration. Laser1: replacing YLF chamber assembly because of flow tube broken.


20-June 05

diCon optical switch 1x80 is going to be shipped to repair --> no more control of the display.


24-May 05

- TiS box3 repair: red LED replaced.

- Particule counter borrow from C. Frei


13-May 05

- Laser1&2 runs. Small tunning of LBO.

- Laser3 TiS box3 not working on Computer Mode but everything still fine on Manual Mode.

Cable OK because TiSbox2 works fine on Laser3. Computer mode it's ok time to time few minutes. Contact Sriram from Quantronix for help.


11May 05

- Run the FFUnits on the 3 laser rooms.

- Reconnection of fibres in each laser (remove for cleanroom installation) with cleanning.


26-April 05

- Softwall crates in H4 (pictures)

- Spare rack: replace CAMAC crate with a new one. Add a NIM crate. Saclay crate free.

- Runs 3 lasers 1 day before installation of softwalls.


25 April 05

Received softwalls at Meyrin site. Internal transport automatic for Prevessin (no EDH). One crate's damaged.


13 April 05

Discovered that CAMAC crate used for Laser1 not working anymore (24 and 12 V). Thus previous plots for Laser1 not reliable.